### v1

# Convergence Rates for Stochastic Proximal and Projection Estimators

Abstract.
In this paper, we establish explicit, non-asymptotic convergence rates for the stochastic smooth approximations of infimal convolutions introduced and developed in [5, 6]. In particular, we quantify the convergence of the associated barycentric estimators toward proximal mappings and metric projections. We prove a dimension-explicit $\sqrt{\delta}$ bound in the $\rho$-weakly convex (possibly nonsmooth) setting and show, by examples, that this order is sharp. Under additional regularity, namely $C^{2}$ smoothness with globally Lipschitz Hessian, we derive an improved linear $O(\delta)$ rate with explicit constants, and we obtain refined projection estimates for convex sets with local $C^{2,1}$ boundary.
Key words and phrases:
proximal mapping, weakly convex functions, stochastic approximation, convergence rates
2020 Mathematics Subject Classification:
Primary 49J53; Secondary 49J42, 90C30, 90C25
1. Introduction
Proximal mappings and metric projections are fundamental primitives in variational analysis and optimization. Given a proper function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ and $\lambda>0$, the proximal point $\operatorname{prox}_{\lambda f}(x)$ is the unique minimizer of $y\mapsto f(y)+\frac{1}{2\lambda}\|x-y\|^{2}$ whenever $f$ is $\rho$-weakly convex and $\lambda\in(0,1/\rho)$. When $f=\iota_{C}$ is the indicator of a closed convex set $C$, the proximal mapping reduces to the metric projection $\operatorname{proj}_{C}(x)$. In many applications, however, evaluating $\operatorname{prox}_{\lambda f}(x)$ or $\operatorname{proj}_{C}(x)$ exactly is computationally expensive, and in black-box settings only function values (possibly noisy) are available.
This paper studies a family of stochastic (zeroth-order) estimators for proximal points and projections based on Gaussian perturbations and exponential reweighting. For $\delta>0$, define the barycenter
(1) $m_{\delta}(x):=\frac{\mathbb{E}_{Y\sim\mathcal{N}(x,\delta\lambda I)}\!\big[Y\,\exp(-f(Y)/\delta)\big]}{\mathbb{E}_{Y\sim\mathcal{N}(x,\delta\lambda I)}\!\big[\exp(-f(Y)/\delta)\big]}.$
When $f=\iota_{C}$, the estimator becomes the conditional mean
$p_{\delta}(x)=\mathbb{E}\big[Y\mid Y\in C\big],\qquad Y\sim\mathcal{N}(x,\delta I),$
which can be interpreted as a “smoothed projection” onto $C$.
The estimators (1) arise naturally from Laplace-type asymptotics for integrals with kernel $\exp(-g/\delta)$, where $g(y):=f(y)+\frac{1}{2\lambda}\|x-y\|^{2}$. Two recent strands of work developed this idea.
First, motivated by Hamilton-Jacobi equations, [5] proposed the HJ-Prox method and proved that, for $\rho$-weakly convex objectives under mild regularity, the barycenter converges to the proximal point as $\delta\to 0^{+}$. Second, [6] places the relationship between infimal convolutions and self-normalized Laplace approximation front and center, extending the construction beyond the quadratic kernel and establishing asymptotic validity under weak local assumptions; in particular, their framework includes smoothed set projections $p_{\delta}(x)=\mathbb{E}[Y\mid Y\in C]$ and proves $p_{\delta(x)}\to\operatorname{proj}_{C}(x)$ as $\delta\to 0^{+}$ under mild local geometric conditions near the projection point.
While these references provide a compelling unifying viewpoint and broad asymptotic consistency results, they leave open a basic quantitative question that is decisive for algorithmic use: how fast do these stochastic estimators converge, and how do the constants depend on the ambient dimension and on the geometry/smoothness of the problem?
Our contributions: explicit and sharp convergence rates
The main goal of this work is to provide non-asymptotic, dimension-explicit convergence rates for (1) and its projection specialization, and to identify regimes where improved rates hold.
(i) Nonsmooth weakly convex case: a sharp $\sqrt{\delta}$ rate. Assume $f$ is proper, lower semicontinuous, and $\rho$-weakly convex, and fix $\lambda\in(0,1/\rho)$ with $\mu:=\frac{1}{\lambda}-\rho>0$. Under the mild geometric condition that $\operatorname{dom}f$ has nonempty interior, we prove the global bound (valid for every $\delta>0$)
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\sqrt{\frac{n\delta}{\mu}}.$
This estimate is complemented by explicit examples showing that the $\sqrt{\delta}$ order cannot be improved in general.
(ii) Smooth case: an $O(\delta)$ rate with explicit constants. If $f\in C^{2}(\mathbb{R}^{n})$ is $\rho$-weakly convex and $\nabla^{2}f$ is globally Lipschitz with constant $L$, we establish the linear-rate refinement
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{3nL}{2\mu^{2}}\,\delta,\qquad\mu=\frac{1}{\lambda}-\rho,$
again for every $\delta>0$. This result quantifies a phenomenon suggested by the smoothing interpretations in [5, 6]: additional curvature regularity improves the bias of the barycentric estimator.
(iii) Projection corollaries and boundary-sensitive refinements. Taking $f=\iota_{C}$ and $\lambda=1$ yields a quantitative convergence estimate for convex projections:
$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|\leq\sqrt{n\,\delta}.$
Beyond the convex Lipschitz regime, we analyze the case of convex sets with $C^{2,1}$ boundary charts at the projection point and obtain a refined $O(\delta)$ expansion. This provides a bridge between the purely asymptotic projection consistency statement in [6] and the boundary-geometry-dependent expansions needed in fine asymptotic analysis.
The remainder of the paper is organized as follows. Section 2 collects the mathematical preliminaries used throughout. Section 3 establishes the $\sqrt{\delta}$ bound for nonsmooth weakly convex functions, provides sharpness examples, and derives the corollary for convex projections. Section 4 proves the $O(\delta)$ rate under $C^{2}$ regularity with Lipschitz continuous Hessian and develops a refined analysis for convex sets with $C^{2,1}$ boundaries. The paper ends with some concluding remarks and possible lines of future research.
2. Mathematical Preliminaries
Throughout this paper, we work in the finite-dimensional Hilbert space $\mathbb{R}^{n}$ endowed with the inner product $\langle\cdot,\cdot\rangle$ and the associated norm $\|\cdot\|$. The closed unit ball in $\mathbb{R}^{n}$ is denoted by $\mathbb{B}$. We use the convention $\exp(+\infty)=+\infty$ and $\exp(-\infty)=0$.
For a ser $A\subset\mathbb{R}^{n}$, the indicator function of $A$ is defined by
$\iota_{A}(x):=\left\{\begin{array}[]{cc}0&\text{ if }x\in A,\\ +\infty&\text{ if }x\notin A.\end{array}\right$.
The distance function from $x\in\mathbb{R}^{n}$ to $A$ is given by
$d_{A}(x)=\inf\left\{\|x-y\|:y\in A\right\}.$
Moreover, if $A$ is nonempty, closed and convex, the projection of $x\in\mathbb{R}^{n}$ onto $A$ is denoted by $\operatorname{proj}_{A}(x)$. Given a function $g\colon\mathbb{R}^{n}\rightarrow\mathbb{\mathbb{R}}\cup\{+\infty\}$, the domain of $g$ is
$\operatorname{dom}(g):=\{x\in\mathbb{R}^{n}:g(x)<+\infty\}.$
We say that $g$ is proper if $\operatorname{dom}g\neq\emptyset$. Given $\rho\geq 0$, a function $g\colon\mathbb{R}^{n}\to\mathbb{\mathbb{R}}\cup\{+\infty\}$ is called $\rho$-weakly convex if $g+\frac{\rho}{2}\|\cdot\|^{2}$ is convex. For $\mu>0$, we say that $g$ is $\mu$-strongly convex if $g-\frac{\mu}{2}\|\cdot\|^{2}$ is convex.
The Moreau envelope of index $\lambda>0$ of a function $g\colon\mathbb{R}^{n}\to{\mathbb{R}}_{\infty}$ is denoted by $\operatorname{e}_{\lambda}g\colon\mathbb{R}^{n}\to\mathbb{R}$ and is defined by
$\operatorname{e}_{\lambda}g(x):=\inf_{y\in\mathbb{R}^{n}}\left(g(y)+\frac{1}{2\lambda}\|x-y\|^{2}\right)\textrm{ for all }x\in\mathbb{R}^{n}.$
If $g$ is proper, lower semicontinuous, and $\rho$-weakly convex, and if $0<\lambda<\frac{1}{\rho}$ (with the convention $1/0=+\infty$), then the above infimum is attained at a unique point $\operatorname{prox}_{\lambda g}(x)\in\mathbb{R}^{n}$, and
$\operatorname{e}_{\lambda}g(x)$ $=g(\operatorname{prox}_{\lambda g}(x))+\frac{1}{2\lambda}\|x-\operatorname{prox}_{\lambda g}(x)\|^{2}.$
In this case, the operator $x\mapsto\operatorname{prox}_{\lambda g}(x)$ is everywhere defined and is called the proximal operator of $g$ of index $\lambda$. Moreover, $\operatorname{e}_{\lambda}g$ is continuously differentiable and, for each $x\in\mathbb{R}^{n}$,
$\nabla\operatorname{e}_{\lambda}g(x)=\frac{1}{\lambda}(x-\operatorname{prox}_{\lambda g}(x)).$
We refer to [1, 2] for more details.
Let $\delta>0$, and let $g\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ be a proper function such that $e^{-g/\delta}$ is integrable and
$\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy>0.$
Under these assumptions, $e^{-g/\delta}$ induces a probability measure. We denote by
$\mathbb{E}_{\sigma_{\delta}}[w(y)]=\frac{\int_{\mathbb{R}^{n}}w(y)e^{-g(y)/\delta}dy}{\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy}$
the expectation of a given function $w$ with respect to this probability measure.
The following result is known as Fradelizi’s inequality (see [4])
Lemma 2.1.
Let $h:\mathbb{R}^{n}\to[0,\infty)$ be an integrable log-concave function, and let
$b_{h}:=\frac{\int_{\mathbb{R}^{n}}y\,h(y)\,dy}{\int_{\mathbb{R}^{n}}h(y)\,dy}$
denotes its barycenter. Then
$h(b_{h})\ \geq\ e^{-n}\,\sup_{z\in\mathbb{R}^{n}}h(z).$
3. Nonsmooth results
Let $f\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ be a $\rho$-weakly convex function with $\rho\geq 0$. The aim of this section is to prove that, for any $0<\lambda<1/\rho$,
$m_{\delta}(x):=\frac{\mathbb{E}_{y\sim\mathcal{N}(x,\delta\lambda I)}[y\cdot\exp(-f(y)/\delta)]}{\mathbb{E}_{y\sim\mathcal{N}(x,\delta\lambda I)}[\exp(-f(y)/\delta)]}\to\operatorname{prox}_{\lambda f}(x)\textrm{ as }\delta\to 0^{+}.$
The next result provides a rate of convergence of the barycenter toward the projection.
Theorem 3.1.
Let $f\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ be proper, lower semicontinuous, and $\rho$-weakly convex for some $\rho\geq 0$. Assume that $\operatorname{dom}f$ has nonempty interior. Fix $x\in\mathbb{R}^{n}$ and $0<\lambda<1/\rho$, and set $\mu:=\frac{1}{\lambda}-\rho>0$. Then, for every $\delta>0$,
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\sqrt{\frac{n\delta}{\mu}}.$
Proof.
Fix $x\in\mathbb{R}^{n}$ and assume that $0<\lambda<1/\rho$. Consider
$g(y):=f(y)+\frac{1}{2\lambda}\|x-y\|^{2}.$
Since $f$ is $\rho$-weakly convex, it follows that $g$ is $\mu$-strongly convex and
(2) $g(y)\geq g(\operatorname{prox}_{\lambda f}(x))+\frac{\mu}{2}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\textrm{ for all }y\in\mathbb{R}^{n}.$
Consider the vector field
$F(y):=(y-\operatorname{prox}_{\lambda f}(x))e^{-g(y)/\delta}.$
Claim 1: $\mathbb{E}_{\sigma_{\delta}}\left[\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle\right]=n\delta$.
Proof of Claim 1: For $R>0$, apply the divergence theorem on $\mathbb{B}_{R}$:
$\int_{\mathbb{B}_{R}}\operatorname{div}F(y)dy=\int_{\partial\mathbb{B}_{R}}\langle F(y),\nu(y)\rangle dS(y).$
Using (2), the right-hand side is bounded in absolutely value by
$CR^{n}\exp\left(-\frac{\mu}{2\delta}(R-\|\operatorname{prox}_{\lambda f}(x)\|)^{2}\right)\to 0\textrm{ as }R\to\infty,$
hence $\int_{\mathbb{R}^{n}}\operatorname{div}F(y)dy=0$. Since $g$ is convex, it is differentiable a.e. and
$\operatorname{div}F(y)dy=ne^{-g(y)/\delta}-\frac{1}{\delta}\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle e^{-g(y)/\delta}\textrm{ for a.e. }y\in\mathbb{R}^{n}.$
Finally, integrating over $\mathbb{R}^{n}$ and dividing by $\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy$, we obtain the claim.
Next, by strong convexity and almost-everywhere differentiability, we obtain
$\langle\nabla g(y),y-\operatorname{prox}_{\lambda f}(x)\rangle\geq\mu\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}.$
Taking expectation yields
$\mu\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\right]\leq\mathbb{E}_{\sigma_{\delta}}\left[\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle\right]=n\delta,$
hence $\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\right]\leq n\delta/\mu$. Finally, by Jensen inequality,
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2}$ $\leq\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|\right]^{2}$
$\leq\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\right]$
$\leq\frac{n\delta}{\mu}.$
Remark 3.2.
Let $Y\sim\sigma_{\delta}$, where $\sigma_{\delta}$ has density proportional to $\exp(-g/\delta)$. The proof of Theorem 3.1 yields the mean-square localization estimate
$\mathbb{E}\big[\|Y-\operatorname{prox}_{\lambda f}(x)\|^{2}\big]\leq\frac{n\delta}{\mu}.$
This bound also yields a simple concentration inequality: for every $r>0$,
$\sigma_{\delta}\!\big(\|Y-\operatorname{prox}_{\lambda f}(x)\|\geq r\big)\leq\frac{n\delta}{\mu\,r^{2}},$
and therefore, for any $\eta\in(0,1)$,
$\sigma_{\delta}\!\left(B_{\sqrt{n\delta/(\mu\eta)}}(\operatorname{prox}_{\lambda f}(x))\right)\geq 1-\eta.$
The following two examples show that our convergence rates are sharp.
Example 3.1.
Fix $n\geq 1$ and consider
$f(y)=\sum_{i=1}^{n}\max\{y_{i},0\}.$
Then $f$ is proper, lower semicontinuous, convex and $\operatorname{dom}f=\mathbb{R}^{n}$ has nonempty interior. Fix $x=0$ and any $\lambda>0$, and set $\mu=1/\lambda$. Then, $\operatorname{prox}_{\lambda f}(0)=0$ and, by symmetry of $f$, for any $\delta>0$
$m_{\delta}(0)=(a_{\delta},\ldots,a_{\delta}),$
where
$a_{\delta}:=\frac{\int_{\mathbb{R}}t\,e^{-\phi(t)/\delta}\,dt}{\int_{\mathbb{R}}e^{-\phi(t)/\delta}\,dt}\qquad\phi(t):=\max\{t,0\}+\frac{1}{2\lambda}t^{2}.$
Moreover, one can show that
$a_{\delta}=-\sqrt{\frac{2}{\pi}\lambda\delta}+O(\delta)\textrm{ as }\delta\to 0^{+}.$
Consequently,
$\|m_{\delta}(0)-\operatorname{prox}_{\lambda f}(0)\|=\sqrt{n}|a_{\delta}|=\sqrt{\frac{2n}{\pi\mu}\delta}+O(\delta).$
This shows that the order of convergence in Theorem 3.1 is sharp.
The following example shows that the bound obtained in Theorem 3.1 is sharp.
Example 3.2.
Fix $n\geq 2$, take $\lambda>0$, set $\mu=1/\lambda$, and let $x=0$. For $\alpha\in(0,\pi/2)$, define the circular cone
$K_{\alpha}:=\{y\in\mathbb{R}^{n}\colon\langle y,e_{1}\rangle\geq\|y\|\,\cos\alpha\}.$
Let $f=\iota_{K_{\alpha}}$ be the indicator function of $K_{\alpha}$. Then $\operatorname{prox}_{\lambda f}(0)=0$. Moreover,
$\|p_{\delta}(0)-\operatorname{prox}_{\lambda f}(0)\|=\|m_{\delta}(0)\|=\sqrt{\frac{\delta}{\mu}}\mathbb{E}[\chi_{n}]m_{\alpha,n},$
where $\chi_{n}$ denotes the $\chi$-distribution with $n$ degrees of freedom and
$m_{\alpha,n}:=\mathbb{E}\left[U_{1}\mid U_{1}\geq\cos\alpha\right],$
with $U$ uniformly distributed on $\mathbb{S}^{n-1}$ and $U_{1}=\langle U,e_{1}\rangle$. In particular, since $U_{1}\in[\cos\alpha,1]$ on the conditioning event, we have $m_{\alpha,n}\in[\cos\alpha,1]$ and hence $m_{\alpha,n}\to 1$ as $\alpha\searrow 0$. Furthermore,
$\mathbb{E}[\chi_{n}]\sqrt{n}\to 1\textrm{ as }n\to\infty.$
Consequently, for every $\varepsilon>0$ there exist $n$ and $\alpha$ such that, for all $\delta>0$,
$\|p_{\delta}(0)-\operatorname{prox}_{\lambda f}(0)\|\geq(1-\varepsilon)\sqrt{\frac{n\delta}{\mu}}.$
This shows that the rate of convergence in Theorem 3.1 is sharp.
We end this section by showing that $m_{\delta}(x)$ coincides with $\operatorname{prox}_{\lambda f}(x)$ whenever $f$ is quadratic.
Example 3.3.
Let $A$ be a symmetric and satisfy $A\succeq-\rho I$ for some $\rho\geq 0$, and define
$f(y)=\frac{1}{2}\langle Ay,y\rangle+\langle b,y\rangle+c.$
Then for every $x\in\mathbb{R}^{n}$, every $\delta>0$, and every $0<\lambda<1/\rho$ (with the convention $1/0=+\infty$), one has
$m_{\delta}(x)=\operatorname{prox}_{\lambda f}(x)=\left(A+\frac{1}{\lambda}I\right)^{-1}\left(\frac{1}{\lambda}x-b\right).$
3.1. Stochastic Approximation of Projections onto Convex Sets
Let $C\subset\mathbb{R}^{n}$ be a nonempty closed convex set with nonempty interior. We define
$p_{\delta}(x):=\frac{\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[y\cdot\mathbf{1}_{C}(y)]}{\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[\mathbf{1}_{C}(y)]}=\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[y\mid y\in C].$
As a direct consequence of Theorem 3.1, we obtain a quantified convergence result: $p_{\delta}(x)\to\operatorname{proj}_{C}(x)$ as $\delta\to 0^{+}$.
Corollary 3.3.
Let $C\subset\mathbb{R}^{n}$ be a nonempty closed convex set with nonempty interior. Fix $x\in\mathbb{R}^{n}$. Then, for every $\delta>0$,
$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|\leq\sqrt{n\,\delta}.$
4. Smooth results
In this section, we show that additional smoothness of the problem data yields improved convergence rates compared with those in Section 3.
Theorem 4.1.
Let $f\colon\mathbb{R}^{n}\to\mathbb{R}$ be $C^{2}$ and $\rho$-weakly convex with $\rho\geq 0$. Assume that $\nabla^{2}f$ is globally Lipschitz: there exists $L\geq 0$ such that
$\|\nabla^{2}f(u)-\nabla^{2}f(v)\|\leq L\|u-v\|\quad\textrm{ for all }u,v\in\mathbb{R}^{n}.$
Fix $x\in\mathbb{R}^{n}$ and $0<\lambda<1/\rho$, and set $\mu:=\frac{1}{\lambda}-\rho>0$. Then, for every $\delta>0$,
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{3nL}{2\mu^{2}}\,\delta.$
Proof.
Fix $x\in\mathbb{R}^{n}$ and assume that $0<\lambda<1/\rho$. Let us consider
$g(y):=f(y)+\frac{1}{2\lambda}\|x-y\|^{2}.$
Since $f$ is $\rho$-weakly convex and $0<\lambda<1/\rho$, it follows that $g$ is $\mu$-strongly convex and
(3) $g(y)\geq g(\operatorname{prox}_{\lambda f}(x))+\frac{\mu}{2}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\textrm{ for all }y\in\mathbb{R}^{n}.$
From this inequality we get that $e^{-g/\delta}$ is integrable for every $\delta>0$, so $m_{\delta}(x)$ is well-defined.
Given a function $w(\cdot)$, we define
$\mathbb{E}_{\sigma_{\delta}}[w(y)]=\frac{\int_{\mathbb{R}^{n}}w(y)e^{-g(y)/\delta}dy}{\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy}$
Claim 1: $\mathbb{E}_{\sigma_{\delta}}[\nabla g(y)]=0$.
Proof of Claim 1: We observe that
$\frac{\partial}{\partial y_{i}}\left(e^{-g(y)/\delta}\right)=-\frac{1}{\delta}\frac{\partial g}{\partial y_{i}}(y)e^{-g(y)/\delta}.$
Moreover, from (3), it follows that
$0\leq e^{-g(y)/\delta}\leq e^{-g(\operatorname{prox}_{\lambda f}(x))/\delta}\cdot e^{-\frac{\mu}{\delta}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}}\textrm{ for all }y\in\mathbb{R}^{n},$
which implies that
$0=\int_{\mathbb{R}^{n}}\frac{\partial}{\partial y_{i}}\left(e^{-g(y)/\delta}\right)dy=-\frac{1}{\delta}\int_{\mathbb{R}^{n}}\frac{\partial g}{\partial y_{i}}(y)e^{-g(y)/\delta}dy,$
which implies the claim.
Claim 2: Let $H=\nabla^{2}g(\operatorname{prox}_{\lambda f}(x))$. Then, for all $y\in\mathbb{R}^{n}$
$\nabla g(y)=H(y-\operatorname{prox}_{\lambda f}(x))+r(y),$
where $\|r(y)\|\leq\frac{L}{2}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}$.
Proof of Claim 2: Since $\nabla g(\operatorname{prox}_{\lambda f}(x))=0$, the formula follows from Taylor’s theorem with
$r(y)=\int_{0}^{1}(\nabla^{2}g(\operatorname{prox}_{\lambda f}(x)+t(y-\operatorname{prox}_{\lambda f}(x)))-H)dt.$
Because $\nabla^{2}g=\nabla^{2}f+\frac{1}{\lambda}I$ and $\nabla^{2}f$ is $L$-Lipschitz, we have
$\|\nabla^{2}g(u)-\nabla^{2}g(v)\|\leq L\|u-v\|.$
Hence
$\|r(y)\|\leq\int_{0}^{1}Lt\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}=\frac{L}{2}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2},$
which proves the claim.
Claim 3: $m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)=-H^{-1}\mathbb{E}_{\sigma_{\delta}}[r(y)]$.
Proof of Claim 3: From Claim 1 and 2 we have that
$0$ $=\mathbb{E}_{\sigma_{\delta}}[\nabla g(y)]=H(m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x))+\mathbb{E}_{\sigma_{\delta}}[r(y)],$
which implies the result.
Claim 4: For every $\delta>0$, we have
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{3nL}{2\mu^{2}}\delta.$
Proof of Claim 4: On the one hand, from Claims 2 and 3, we have
(4) $\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|$ $=\|H^{-1}\mathbb{E}_{\sigma_{\delta}}[r(y)]\|$
$\leq\frac{1}{\mu}\mathbb{E}_{\sigma_{\delta}}\left[\|r(y)\|\right]$
$\leq\frac{L}{2\mu}\mathbb{E}_{\sigma_{\delta}}[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}],$
where we have used that $H\succcurlyeq\mu I$ and $\|H^{-1}\|\leq 1/\mu$. Moreover,
(5) $\mathbb{E}_{\sigma_{\delta}}[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}]=\mathbb{E}_{\sigma_{\delta}}[\|y-m_{\delta}(x)\|^{2}]+\mathbb{E}_{\sigma_{\delta}}[\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2}].$
On the other hand, we observe that $g/\delta$ is $\mu/\delta$ strongly convex. Hence, $h_{\delta}=e^{-g/\delta}$ is strongly log-concave. From Brascamp–Lieb inequality (see [3, Theorem 4.1]), we get that
(6) $\operatorname{Cov}_{\sigma_{\delta}}(y)\preceq\frac{\delta}{\mu}I\quad\Rightarrow\quad\mathbb{E}_{\sigma_{\delta}}[\|y-m_{\delta}(x)\|^{2}]=\operatorname{tr}\left(\operatorname{Cov}_{\sigma_{\delta}}(y)\right)\leq\operatorname{tr}\left(\frac{\delta}{\mu}I\right)=\frac{n\delta}{\mu}.$
Next, again since $h_{\delta}=e^{-g/\delta}$ is log-concave, applying the Lemma 2.1 to $h_{\delta}$ yields
(7) $g(m_{\delta}(x))-g(\operatorname{prox}_{\lambda f}(x))\leq n\delta.$
By $\mu$-strong convexity of $g$,
$g(m_{\delta}(x))-g(\operatorname{prox}_{\lambda f}(x))\geq\frac{\mu}{2}\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2},$
and combining with (7) gives
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2}\leq\frac{2n\delta}{\mu}.$
Plugging (6) and the above inequality into (5) yields
$\mathbb{E}_{\sigma_{\delta}}[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}]\leq\frac{n\delta}{\mu}+\frac{2n\delta}{\mu}=\frac{3n\delta}{\mu}.$
Finally, from the above inequality and (4), we get
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{L}{2\mu}\cdot\frac{3n\delta}{\mu}=\frac{3nL}{2\mu^{2}}\delta,$
which proves Claim 4 and the theorem. ∎
4.1. Stochastic Approximation of Projections onto Smooth Convex Sets
Let $C\subset\mathbb{R}^{n}$ be a nonempty closed convex set with nonempty interior. Recall that
$p_{\delta}(x):=\frac{\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[y\cdot\mathbf{1}_{C}(y)]}{\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[\mathbf{1}_{C}(y)]}=\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[y\mid y\in C].$
Definition 4.2.
Let $C\subset\mathbb{R}^{n}$ be a closed set and let $p\in\partial C$. We say that $\partial C$ admits a local $C^{2,1}$ chart at $p$ with constants $(\rho,L,M)$ if, after a rigid motion sending $p$ to 0 and the outward unit normal at $p$ to $e_{n}$, there exists $\rho>0$ and a concave function $h\in C^{2,1}(\mathbb{B}^{n-1}_{\rho})$ such that $h(0)=0$, $\nabla h(0)=0$, $\|\nabla^{2}h(z)\|\leq L$,
$C\cap\big(\mathbb{B}^{n-1}_{\rho}\times(-\rho,\rho)\big)=\{(z,t)\in\mathbb{R}^{n-1}\times\mathbb{R}:\ t\leq h(z)\},$
and $\nabla^{2}h$ is Lipschitz on $\mathbb{B}^{n-1}_{\rho}$ with constant $M\geq 0$, i.e.,
$\|\nabla^{2}h(z)-\nabla^{2}h(z^{\prime})\|\leq M\|z-z^{\prime}\|\qquad\text{for all }z,z^{\prime}\in\mathbb{B}^{n-1}_{\rho}.$
In the case of a smooth convex set, we obtain the following improvement of Corollary 3.3, which provides the convergence rate
$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|=O(\delta)\textrm{ as }\delta\to 0^{+}.$
Proposition 4.3.
Let $C\subset\mathbb{R}^{n}$ be a closed convex set with nonempty interior, and fix $x\in\mathbb{R}^{n}\setminus C$. Assume that $\partial C$ admits a local $C^{2,1}$ chart at $\operatorname{proj}_{C}(x)$ with constants $(\rho,L,M)$. Then there exists $\delta_{0}>0$ such for every $0<\delta<\delta_{0}$, we have
$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|\leq$ $K_{n}\delta,$
where $K_{n}$ is a constant depending on the dimension $n$, $\rho$, $K$, and $M$.
Example 4.1.
Consider the half-space $C=\{y\in\mathbb{R}^{n}:\langle y,n\rangle\leq 0$, where $\|n\|=1$. Then $\partial C=\{y\in\mathbb{R}^{n}:\langle y,n\rangle=0\}$ is a hyperplane. Pick $x\notin C$, so that $\langle x,n\rangle>0$. Then
$\operatorname{proj}_{C}(x)=x-\langle x,n\rangle n,\quad d_{C}(x)=\langle x,n\rangle.$
Moreover, since $p_{\delta}(x)=\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}\!\left[\,y\mid y\in C\,\right]$, symmetry implies that the tangential conditional mean is unchanged and the drift is purely normal. Hence,
$p_{\delta}(x)-\operatorname{proj}_{C}(x)=\alpha_{\delta}\,n$
for some scalar $\alpha_{\delta}<0$. In fact, one can show that
$\alpha_{\delta}:=d_{C}(x)-\sqrt{\delta}\frac{\phi(d_{C}(x)/\sqrt{\delta})}{\Phi(-d_{C}(x)/\sqrt{\delta})},$
and therefore
$p_{\delta}(x)=\operatorname{proj}_{C}(x)+\left(d_{C}(x)-\sqrt{\delta}\frac{\phi(d_{C}(x)/\sqrt{\delta})}{\Phi(-d_{C}(x)/\sqrt{\delta})}\right)n,$
which implies that $\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|=O(\delta)$ as $\delta\to 0^{+}$. Hence, the order of the bound obtained in Proposition 4.3 is sharp.

### v2

@@
 Abstract.
-In this paper, we establish explicit, non-asymptotic convergence rates for the stochastic smooth approximations of infimal convolutions introduced and developed in [5, 6].
+In this paper, we establish explicit convergence rates for the stochastic smooth approximations of infimal convolutions introduced and developed in [9, 10].
 In particular, we quantify the convergence of the associated barycentric estimators toward proximal mappings and metric projections.
-We prove a dimension-explicit $\sqrt{\delta}$ bound in the $\rho$-weakly convex (possibly nonsmooth) setting and show, by examples, that this order is sharp.
+We prove a dimension-explicit $\sqrt{\delta}$ bound, with explicit constants for the proximal mapping, in the $\rho$-weakly convex (possibly nonsmooth) setting, and we also obtain a dimension-explicit $\sqrt{\delta}$ rate for the metric projection onto an arbitrary convex set with nonempty interior.
 Under additional regularity, namely $C^{2}$ smoothness with globally Lipschitz Hessian, we derive an improved linear $O(\delta)$ rate with explicit constants, and we obtain refined projection estimates for convex sets with local $C^{2,1}$ boundary.
+Examples demonstrate that these rates are optimal.
 Key words and phrases:
@@
 Lemma 2.1.
-Let $h:\mathbb{R}^{n}\to[0,\infty)$ be an integrable log-concave function, and let
-$b_{h}:=\frac{\int_{\mathbb{R}^{n}}y\,h(y)\,dy}{\int_{\mathbb{R}^{n}}h(y)\,dy}$
-denotes its barycenter. Then
+Let $h:\mathbb{R}^{n}\to[0,\infty)$ be a log-concave function, not identically zero, such that $\int_{\mathbb{R}^{n}}h(y)\,dy<\infty$. Define its barycenter by
+$b_{h}:=\frac{\int_{\mathbb{R}^{n}}y\,h(y)\,dy}{\int_{\mathbb{R}^{n}}h(y)\,dy}.$
+Then $b_{h}$ is well-defined and
 $h(b_{h})\ \geq\ e^{-n}\,\sup_{z\in\mathbb{R}^{n}}h(z).$
@@
 $F(y):=(y-\operatorname{prox}_{\lambda f}(x))e^{-g(y)/\delta}.$
-Claim 1: $\mathbb{E}_{\sigma_{\delta}}\left[\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle\right]=n\delta$.
-Proof of Claim 1: For $R>0$, apply the divergence theorem on $\mathbb{B}_{R}$:
-$\int_{\mathbb{B}_{R}}\operatorname{div}F(y)dy=\int_{\partial\mathbb{B}_{R}}\langle F(y),\nu(y)\rangle dS(y).$
-Using (2), the right-hand side is bounded in absolutely value by
-$CR^{n}\exp\left(-\frac{\mu}{2\delta}(R-\|\operatorname{prox}_{\lambda f}(x)\|)^{2}\right)\to 0\textrm{ as }R\to\infty,$
-hence $\int_{\mathbb{R}^{n}}\operatorname{div}F(y)dy=0$. Since $g$ is convex, it is differentiable a.e. and
-$\operatorname{div}F(y)dy=ne^{-g(y)/\delta}-\frac{1}{\delta}\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle e^{-g(y)/\delta}\textrm{ for a.e. }y\in\mathbb{R}^{n}.$
-Finally, integrating over $\mathbb{R}^{n}$ and dividing by $\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy$, we obtain the claim.
+Claim 1: $\mathbb{E}_{\sigma_{\delta}}\left[\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle\right]\leq n\delta$.
+Proof of Claim 1: Since $y\mapsto g(y)$ is proper, lsc, and convex, its domain is convex; in particular, $D:=\operatorname{dom}f=\operatorname{dom}g$ is convex with nonempty interior (by assumption).
+On $\operatorname{int}(D)$, every proper lsc convex function is locally Lipschitz, hence differentiable a.e.
+and belongs to $W^{1,1}_{\operatorname{loc}}(\operatorname{int}(D))$.
+Consequently, $\nabla g\in L^{1}_{\operatorname{loc}}(\operatorname{int}(D);\mathbb{R}^{n})$, and for a.e.
+$y\in\operatorname{int}(D)$,
+$\operatorname{div}F(y)=ne^{-g(y)/\delta}-\frac{1}{\delta}\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle e^{-g(y)/\delta}.$
+Fix $R>0$ and set $\Omega_{R}:=\operatorname{int}(D)\cap\mathbb{B}_{R}$. Since
+$\|y-\operatorname{prox}_{\lambda f}(x)\|\leq R+\|\operatorname{prox}_{\lambda f}(x)\|=:C_{R}\textrm{ on }\mathbb{B}_{R}$
+and $e^{-g(y)/\delta}\leq e^{-\inf_{\mathbb{B}_{R}}g/\delta}=:M_{R,\delta}<\infty$, it follows that for a.e. $y\in\Omega_{R}$,
+$|\operatorname{div}F(y)|\leq nM_{R,\delta}+\frac{M_{R,\delta}C_{R}}{\delta}\|\nabla g(y)\|,$
+Since $\nabla g\in L^{1}_{\operatorname{loc}}(\operatorname{int}(D);\mathbb{R}^{n})$, we have $\nabla g\in L^{1}(\Omega_{R};\mathbb{R}^{n})$, and then $\operatorname{div}F\in L^{1}(\Omega_{R})$.
+For $R>0$, apply the divergence theorem on $\Omega_{R}$:
+$\int_{\Omega_{R}}\operatorname{div}F(y)dy$ $=\int_{\partial\Omega_{R}}F\cdot\nu\,d\mathcal{H}^{n-1}.$
+Moreover, by Lemma A.1, we can select $R_{k}\to\infty$ such that
+$\int_{\partial\Omega_{R_{k}}}\langle F,\nu\rangle\,d\mathcal{H}^{n-1}=\underbrace{\int_{\partial D\cap\mathbb{B}_{R_{k}}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}}_{=:I_{1}}+\underbrace{\int_{\partial\mathbb{B}_{R_{k}}\cap D}F\cdot\nu_{\mathbb{B}_{R_{k}}}\,d\mathcal{H}^{n-1}}_{=:I_{2}}.$
+Now, on the one hand, since $D$ is convex and $\operatorname{prox}_{\lambda f}(x)\in D$, for $\mathcal{H}^{n-1}$-a.e. $y\in\partial D$ one has
+$\langle y-\operatorname{prox}_{\lambda f}(x),\nu_{D}(y)\rangle\geq 0\quad\Rightarrow\quad\langle F(y),\nu_{D}(y)\rangle\geq 0.$
+Hence the $\partial D$-term $I_{1}$ is nonnegative. On the other hand, using (2), the term $I_{1}$ is bounded in absolutely value by
+$CR_{k}^{n}\exp\left(-\frac{\mu}{2\delta}(R_{k}-\|\operatorname{prox}_{\lambda f}(x)\|)^{2}\right)\to 0\textrm{ as }k\to\infty,$
+hence the $\partial\mathbb{B}_{R_{k}}$-term $I_{2}$ vanishes as $k\to\infty$. Therefore,
+$0$ $\leq\limsup_{k\to\infty}\int_{\mathbb{B}_{R_{k}}\cap\partial D}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}$
+$=\limsup_{k\to\infty}\Big(\int_{\partial D\cap\mathbb{B}_{R_{k}}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\partial\mathbb{B}_{R_{k}}\cap D}F\cdot\nu_{\mathbb{B}_{R_{k}}}\,d\mathcal{H}^{n-1}\Big)$
+$=\limsup_{k\to\infty}\int_{\partial\Omega_{R_{k}}}F\,\nu\,d\mathcal{H}^{n-1}$
+$=\limsup_{k\to\infty}\int_{\Omega_{R_{k}}}\operatorname{div}F(y)dy$
+$=\int_{D}\operatorname{div}F(y)dy$
+$=n\int_{D}e^{-g(y)/\delta}dy-\frac{1}{\delta}\int_{D}\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle e^{-g(y)/\delta}dy.$
+Finally, dividing by $\int_{D}e^{-g(y)/\delta}dy$ and rearranging, we obtain the claim.∎
 Next, by strong convexity and almost-everywhere differentiability, we obtain
@@
 which proves Claim 4 and the theorem. ∎
-4.1. Stochastic Approximation of Projections onto Smooth Convex Sets
-Let $C\subset\mathbb{R}^{n}$ be a nonempty closed convex set with nonempty interior. Recall that
+4.1. Stochastic Approximation of Projections onto Smooth Sets
+Let $C\subset\mathbb{R}^{n}$ be a nonempty closed set with nonempty interior. Recall that
 $p_{\delta}(x):=\frac{\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[y\cdot\mathbf{1}_{C}(y)]}{\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[\mathbf{1}_{C}(y)]}=\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[y\mid y\in C].$
 Definition 4.2.
-Let $C\subset\mathbb{R}^{n}$ be a closed set and let $p\in\partial C$.
+Let $C\subset\mathbb{R}^{n}$ be a closed convex set and let $p\in\partial C$.
 We say that $\partial C$ admits a local $C^{2,1}$ chart at $p$ with constants $(\rho,L,M)$ if, after a rigid motion sending $p$ to 0 and the outward unit normal at $p$ to $e_{n}$, there exists $\rho>0$ and a concave function $h\in C^{2,1}(\mathbb{B}^{n-1}_{\rho})$ such that $h(0)=0$, $\nabla h(0)=0$, $\|\nabla^{2}h(z)\|\leq L$,
@@
 $\|\nabla^{2}h(z)-\nabla^{2}h(z^{\prime})\|\leq M\|z-z^{\prime}\|\qquad\text{for all }z,z^{\prime}\in\mathbb{B}^{n-1}_{\rho}.$
+Figure 1. Local $C^{2,1}$ chart of $\partial C$.
 In the case of a smooth convex set, we obtain the following improvement of Corollary 3.3, which provides the convergence rate
 $\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|=O(\delta)\textrm{ as }\delta\to 0^{+}.$
-Proposition 4.3.
+Theorem 4.3.
 Let $C\subset\mathbb{R}^{n}$ be a closed convex set with nonempty interior, and fix $x\in\mathbb{R}^{n}\setminus C$.
 Assume that $\partial C$ admits a local $C^{2,1}$ chart at $\operatorname{proj}_{C}(x)$ with constants $(\rho,L,M)$.
-Then there exists $\delta_{0}>0$ such for every $0<\delta<\delta_{0}$, we have
-$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|\leq$ $K_{n}\delta,$
-where $K_{n}$ is a constant depending on the dimension $n$, $\rho$, $K$, and $M$.
+Then there exist $\mathcal{C}>0$ and $\delta_{0}>0$ such for every $0<\delta<\delta_{0}$, we have
+$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|\leq$ $\,\mathcal{C}\delta+\mathcal{O}(\delta^{2}).$
+The next example shows that the order of the bound obtained in Theorem 4.3 is sharp.
 Example 4.1.
-Consider the half-space $C=\{y\in\mathbb{R}^{n}:\langle y,n\rangle\leq 0$, where $\|n\|=1$.
-Then $\partial C=\{y\in\mathbb{R}^{n}:\langle y,n\rangle=0\}$ is a hyperplane.
-Pick $x\notin C$, so that $\langle x,n\rangle>0$.
+Consider the half-space $C=\{y\in\mathbb{R}^{n}:\langle y,n\rangle\leq 0\}$, where $\|\nu\|=1$.
+Then $\partial C=\{y\in\mathbb{R}^{n}:\langle y,\nu\rangle=0\}$ is a hyperplane.
+Pick $x\notin C$, so that $\langle x,\nu\rangle>0$.
 Then
-$\operatorname{proj}_{C}(x)=x-\langle x,n\rangle n,\quad d_{C}(x)=\langle x,n\rangle.$
+$\operatorname{proj}_{C}(x)=x-\langle x,\nu\rangle\nu,\quad d_{C}(x)=\langle x,\nu\rangle.$
 Moreover, since $p_{\delta}(x)=\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}\!\left[\,y\mid y\in C\,\right]$, symmetry implies that the tangential conditional mean is unchanged and the drift is purely normal.
 Hence,
-$p_{\delta}(x)-\operatorname{proj}_{C}(x)=\alpha_{\delta}\,n$
+$p_{\delta}(x)-\operatorname{proj}_{C}(x)=\alpha_{\delta}\,\nu$
 for some scalar $\alpha_{\delta}<0$. In fact, one can show that
@@
 and therefore
-$p_{\delta}(x)=\operatorname{proj}_{C}(x)+\left(d_{C}(x)-\sqrt{\delta}\frac{\phi(d_{C}(x)/\sqrt{\delta})}{\Phi(-d_{C}(x)/\sqrt{\delta})}\right)n,$
-which implies that $\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|=O(\delta)$ as $\delta\to 0^{+}$. Hence, the order of the bound obtained in Proposition 4.3 is sharp.
+$p_{\delta}(x)=\operatorname{proj}_{C}(x)+\left(d_{C}(x)-\sqrt{\delta}\frac{\phi(d_{C}(x)/\sqrt{\delta})}{\Phi(-d_{C}(x)/\sqrt{\delta})}\right)\nu,$
+which implies that $\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|=O(\delta)$ as $\delta\to 0^{+}$. Hence, the order of the bound obtained in Theorem 4.3 is sharp.
+Appendix A Auxiliary lemmas
+Lemma A.1.
+Let $D\subset\mathbb{R}^{n}$ be a closed convex set with nonempty interior, and for $R>0$ set $\Omega_{R}:=\operatorname{int}(D)\cap\mathbb{B}_{R}$.
+Then $\Omega_{R}$ is a bounded Lipschitz domain.
+Moreover, for a.e.
+$R>0$,
+(8) $\mathcal{H}^{n-1}\big(\partial D\cap\partial\mathbb{B}_{R}\big)=0.$
+Let $F\in W^{1,1}_{\operatorname{loc}}(\operatorname{int}(D);\mathbb{R}^{n})$. Then, for every $R>0$ such that (8) holds and $\int_{\Omega_{R}}(|F(x)|+\|\nabla F(x)\|)\,dx<\infty$, we have
+$\int_{\Omega_{R}}\operatorname{div}F(x)\,dx=\int_{\partial D\cap\mathbb{B}_{R}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\operatorname{int}(D)\cap\partial\mathbb{B}_{R}}F\cdot\nu_{\mathbb{B}_{R}}\,d\mathcal{H}^{n-1},$
+where $\nu_{D}$ is the outer unit normal to $D$, defined $\mathcal{H}^{n-1}$-a.e.
+on $\partial D$, $\nu_{\mathbb{B}_{R}}$ is the outer unit normal to $\mathbb{B}_{R}$.
+Here, $F$ on $\partial\Omega_{R}$ denotes the interior Sobolev trace of $F|_{\Omega_{R}}$ on $\partial\Omega_{R}$, restricted to $\partial D\cap\mathbb{B}_{R}$ and to $\operatorname{int}(D)\cap\partial\mathbb{B}_{R}$.
+Proof.
+The first assertion follows from [7, Chapter 1], since every convex domain has a Lipschitz boundary.
+To prove (8), we observe that, since $D$ is convex, its boundary $\partial D$ is countably (n1)(n-1)-rectifiable.
+Hence, there exist countably many Lipschitz maps $\Phi_{k}\colon U_{k}\subset\mathbb{R}^{n-1}\to\mathbb{R}^{n}$ such that
+$\mathcal{H}^{n-1}\left(\partial D\setminus\bigcup_{k=1}^{\infty}\Phi_{k}(U_{k})\right)=0.$
+see, e.g., [5].
+Fix $k$.
+Define the Lipschitz function $g_{k}\colon U_{k}\to\mathbb{R}$ by $g_{k}(u):=\|\Phi_{k}(u)\|$.
+By the coarea formula for Lipschitz functions (see [4, Chapter 1]), for a.e.
+$R>0$ the level set $g_{k}^{-1}(R)$ is (n2)(n-2)-rectifiable and has locally finite $\mathcal{H}^{n-2}$-measure.
+In particular, $\mathcal{L}^{n-1}(g_{k}^{-1}(R))=0$.
+Then, using the area estimate for Lipschitz maps (a consequence of the area formula),
+$\mathcal{H}^{n-1}(\Phi_{k}(g_{k}^{-1}(R)))\leq\operatorname{Lip}(\Phi_{k})^{n-1}\mathcal{L}^{n-1}(g_{k}^{-1}(R))=0.$
+Moreover, $\Phi_{k}(g_{k}^{-1}(R))\subset\partial D\cap\partial\mathbb{B}_{R}$. Hence, for a.e. $R>0$,
+$\mathcal{H}^{n-1}\big((\partial D\cap\partial\mathbb{B}_{R})\cap\Phi_{k}(U_{k})\big)=0.$
+Taking the countable union over $k$ and using the $\mathcal{H}^{n-1}$-negligibility of the uncovered part of $\partial D$, we obtain (8).
+To prove the last assertion, since $\Omega_{R}\subset\operatorname{int}(D)$ is open and bounded, the assumptions on $F$ imply that $F\in W^{1,1}(\Omega_{R};\mathbb{R}^{n})$ and $\operatorname{div}F\in L^{1}(\Omega_{R})$.
+Since $\Omega_{R}$ is a bounded Lipschitz domain, the (interior) trace operator $\operatorname{Tr}\colon W^{1,1}(\Omega_{R})\to L^{1}(\partial\Omega_{R})$ is well-defined, and the Gauss-Green formula for Sobolev vector fields (see, e.g., [8, Chapter 6]) yields
+(9) $\int_{\Omega_{R}}\operatorname{div}F\,dx=\int_{\partial\Omega_{R}}\operatorname{Tr}(F)\cdot\nu_{\Omega_{R}}\,d\mathcal{H}^{n-1}.$
+Next, we observe that
+$\partial\Omega_{R}=(\partial D\cap\mathbb{B}_{R})\cup(\operatorname{int}(D)\cap\partial\mathbb{B}_{R})\cup(\partial D\cap\partial\mathbb{B}_{R}).$
+Moreover, on $\partial D\cap\mathbb{B}_{R}$ the boundary $\partial\Omega_{R}$ coincides locally with $\partial D$.
+Hence $\nu_{\Omega_{R}}=\nu_{D}$ $\mathcal{H}^{n-1}$-a.e.
+on $\partial D\cap\mathbb{B}_{R}$.
+Finally, by (8), the set $\partial D\cap\partial\mathbb{B}_{R}$ has $\mathcal{H}^{n-1}$-measure zero and does not contribute to the boundary integral in (9).
+Therefore (9) splits into
+$\int_{\Omega_{R}}\operatorname{div}F\,dx=\int_{\partial D\cap\mathbb{B}_{R}}\operatorname{Tr}(F)\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\operatorname{int}(D)\cap\partial\mathbb{B}_{R}}\operatorname{Tr}(F)\cdot\nu_{\mathbb{B}_{R}}\,d\mathcal{H}^{n-1},$
+as claimed. ∎
+Lemma A.2.
+Let $n\geq 1$, $R>0$, $d>0$, and $k\in\{0,1\}$. Then, for every $0<\delta\leq\frac{R^{2}}{2n}$, one has
+$\int_{\{y\in\mathbb{R}^{n}:\ \|y\|\geq R\}}(\|y\|+d)^{k}\exp\!\Big(-\frac{\|y\|^{2}}{2\delta}\Big)\,dy\leq\frac{4\pi^{n/2}}{\Gamma\!\left(\frac{n}{2}\right)}(R+d)^{k}R^{n-2}\delta\,\exp\!\Big(-\frac{R^{2}}{2\delta}\Big).$
+Lemma A.3.
+For all $\delta>0$ and $y>0$,
+$\int_{y}^{\infty}e^{-\frac{t^{2}}{2\delta}}\,dt\geq\frac{\delta y}{y^{2}+\delta}e^{-\frac{y^{2}}{2\delta}}.$
+Appendix B Proof of Theorem 4.3
+Step 1: Initialization
+Set $p:=\operatorname{proj}_{C}(x)$, $d:=d_{C}(x)$, and $\nu:=\frac{x-p}{d}$.
+After translating and rotating, we may assume that $p=0$, $\nu=e_{n}$, and $x=de_{n}$.
+For $y\in\mathbb{R}^{n}$, write $y=(z,t)\in\mathbb{R}^{n-1}\times\mathbb{R}$.
+Let $H=\nabla^{2}h(0)$ (hence $H\preceq 0$ by concavity, and $\|H\|\leq L$).
+By Taylor’s theorem with Lipschitz Hessian, for all $z\in\mathbb{B}_{\rho}^{n-1}$.
+(10) $h(z)=\frac{1}{2}\langle Hz,z\rangle+r(z),\quad|r(z)|\leq\frac{M}{6}\|z\|^{3},$
+and also
+(11) $|h(z)|\leq\frac{L}{2}\|z\|^{2}\textrm{ for all }z\in\mathbb{B}_{\rho}^{n-1}.$
+Moreover, by convexity of $C$ and the fact that $e_{n}$ is an outward normal at 0 imply that the supporting hyperplane is $\{t=0\}$, hence $h(z)\leq 0$ on $\mathbb{B}_{\rho}^{n-1}$.
+Define (see Figure 2)
+$C_{\operatorname{loc}}:=C\cap\big(\mathbb{B}^{n-1}_{\rho}\times(-\rho,\rho)\big)\textrm{ and }C_{\operatorname{tail}}:=C\setminus C_{\operatorname{loc}}.$
+Let us define
+$N_{\operatorname{loc}}$ $=\int_{C_{\operatorname{loc}}}ye^{-\frac{\|y-x\|^{2}}{2\delta}}\,dy,$ $N_{\operatorname{tail}}$ $=\int_{C_{\operatorname{tail}}}ye^{-\frac{\|y-x\|^{2}}{2\delta}}\,dy,$
+$D_{\operatorname{loc}}$ $=\int_{C_{\operatorname{loc}}}e^{-\frac{\|y-x\|^{2}}{2\delta}}\,dy,$ $D_{\operatorname{tail}}$ $=\int_{C_{\operatorname{tail}}}e^{-\frac{\|y-x\|^{2}}{2\delta}}\,dy.$
+With the above definitions, we observe that
+$p_{\delta}(x):=\frac{N_{\operatorname{loc}}+N_{\operatorname{tail}}}{D_{\operatorname{loc}}+D_{\operatorname{tail}}}.$
+Step 2: Bounds on $D_{\operatorname{tail}}$ and $D_{\operatorname{loc}}$.
+We observe that $C_{\operatorname{loc}}\subset\mathbb{B}_{\rho}^{n-1}\times(-\rho,0]$ and if $y\in C_{\operatorname{tail}}$, then either $\|z\|\geq\rho$ or $t\leq-\rho$. In both cases,
+(12) $\|y-x\|^{2}=\|z\|^{2}+(t-d)^{2}\geq\rho^{2}+d^{2}\textrm{ for }y\in C_{\operatorname{tail}}.$
+Figure 2. Local chart of $\partial C$ and the sets $C_{\operatorname{loc}}$ (pale turquoise) and $C_{\operatorname{tail}}$ (light cyan).
+Claim 1: Assume that $0<\delta\leq\frac{\rho^{2}+d^{2}}{2n}$ (with the convention that $1/0=+\infty$). Then
+$D_{\operatorname{tail}}:=\int_{C_{\operatorname{tail}}}e^{-\frac{\|y-x\|^{2}}{2\delta}}dy\leq\frac{4\pi^{n/2}}{\Gamma\left(\frac{n}{2}\right)}(\rho^{2}+d^{2})^{\frac{n-2}{2}}\delta e^{-\frac{\rho^{2}+d^{2}}{2\delta}}.$
+Proof of Claim 1: Since $C_{\operatorname{tail}}\subset\{y\in\mathbb{R}^{n}:\|y-x\|^{2}\geq\rho^{2}+d^{2}\}$, we have
+$D_{\operatorname{tail}}=\int_{C_{\operatorname{tail}}}e^{-\frac{\|y-x\|^{2}}{2\delta}}dy\leq\int_{\{y\in\mathbb{R}^{n}:\|y-x\|^{2}\geq\rho^{2}+d^{2}\}}e^{-\frac{\|y-x\|^{2}}{2\delta}}dy.$
+Applying Lemma A.2 with $R=\sqrt{\rho^{2}+d^{2}}$ yields the claim. ∎
+Claim 2: Assume that $0<\delta\leq\min\left\{\rho^{2}(1+dL),\frac{2\rho(1+dL)}{L},\frac{d\rho}{1+\frac{dL}{2(1+dL)}}\right\}$.
+Then
+$D_{\operatorname{loc}}:=\int_{C_{\operatorname{loc}}}e^{-\frac{\|y-x\|^{2}}{2\delta}}dy\geq\left[\psi(\delta)\cdot\mathcal{K}_{\operatorname{err}}(\delta)\cdot\frac{C_{\operatorname{vol}}(n)}{(1+dL)^{\frac{n-1}{2}}}\right]\cdot\delta^{\frac{n-1}{2}}\cdot e^{-\frac{d^{2}}{2\delta}},$
+where
+$\psi(\delta)$ $=\inf_{z\in\mathbb{B}^{n-1}_{r_{\delta}}}\omega(z),\quad r_{\delta}=\sqrt{\frac{\delta}{1+dL}},$
+$\omega(z)$ $=\frac{\delta(d-h(z))}{(d-h(z))^{2}+\delta}-\frac{\delta}{\rho+d}\exp\left(-\frac{\left[(\rho+d)^{2}-(d-h(z))^{2}\right]}{2\delta}\right),$
+$\mathcal{K}_{\operatorname{err}}(\delta)$ $=\exp\left(-\frac{L^{2}\delta}{8(1+dL)^{2}}\right),$
+$C_{\operatorname{vol}}(n)$ $=\int_{\mathbb{B}^{n-1}}e^{-\frac{\|v\|^{2}}{2}}dv.$
+Proof of Claim 2: Since $\delta\leq\rho^{2}(1+dL)$, we have that $r_{\delta}\leq\rho$. Set A(z):=dh(z)A(z):=d-h(z) and observe that
+  • By (11), we have
+    $A(z)=d-h(z)\leq d+\frac{L}{2}r_{\delta}^{2}=d+\frac{L\delta}{2(1+dL)}=:A_{\max}.$
+  • Since $0<\delta<\frac{2\rho(1+dL)}{L}$, it follows that $A_{\max}<d+\rho$. Consequently,
+    $(d+\rho)^{2}-A(z)^{2}\geq(d+\rho)^{2}-A_{\max}^{2}.$
+  • Since $0<\delta<\frac{d\rho}{1+\frac{dL}{2(1+dL)}}$, we obtain the lower bound
+    $J(z)$ $=\int_{A(z)}^{d+\rho}e^{-\frac{t^{2}}{2\delta}}\,dt$
+    $=\int_{A(z)}^{+\infty}e^{-\frac{t^{2}}{2\delta}}\,dt-\int_{\rho+d}^{\infty}e^{-\frac{t^{2}}{2\delta}}\,dt$
+    $\geq\frac{\delta A(z)}{A(z)^{2}+\delta}e^{-\frac{A(z)^{2}}{2\delta}}-\frac{\delta}{\rho+d}e^{-\frac{(\rho+d)^{2}}{2\delta}}$
+    $=e^{-\frac{A(z)^{2}}{2\delta}}\omega(z)$
+    $\geq\psi(\delta)\,e^{-\frac{A(z)^{2}}{2\delta}},$
+    where we have used Lemma A.3.
+Therefore,
+$D_{\operatorname{loc}}$ $=\int_{z\in\mathbb{B}_{\rho}^{n-1}}e^{-\frac{\|z\|^{2}}{2\delta}}\left(\int_{-\rho}^{h(z)}e^{-\frac{(t-d)^{2}}{2\delta}}\,dt\right)\,dz$
+$=\int_{z\in\mathbb{B}_{\rho}^{n-1}}e^{-\frac{\|z\|^{2}}{2\delta}}J(z)\,dz$
+$\geq\int_{z\in\mathbb{B}_{r_{\delta}}^{n-1}}\psi(\delta)e^{-\frac{\|z\|^{2}+A(z)^{2}}{2\delta}}\,dz.$
+Define $E(z):=\|z\|^{2}+A(z)^{2}$. Since $\|z\|\leq r_{\delta}$, it follows that
+$E(z)\leq d^{2}+(1+dL)\|z\|^{2}+\frac{L^{2}}{4}\|z\|^{4}\leq d^{2}+(1+dL)\|z\|^{2}+\frac{L^{2}}{4}\left(\frac{\delta}{1+dL}\right)^{2}.$
+Hence,
+$D_{\operatorname{loc}}$ $\geq\psi(\delta)e^{-\frac{L^{2}\delta}{8(1+dL)^{2}}}e^{-\frac{d^{2}}{2\delta}}\left(\frac{\delta}{1+dL}\right)^{\frac{n-1}{2}}\int_{\mathbb{B}^{n-1}}e^{-\frac{\|v\|^{2}}{2}}dv$
+$=\left[\psi(\delta)\cdot\mathcal{K}_{\operatorname{err}}(\delta)\cdot\frac{C_{\operatorname{vol}}(n)}{(1+dL)^{\frac{n-1}{2}}}\right]\cdot\delta^{\frac{n-1}{2}}\cdot e^{-\frac{d^{2}}{2\delta}},$
+which proves the claim.
+Claim 3: If $0<\delta\leq\delta_{0}:=\min\big\{\rho^{2}(1+dL),\frac{2\rho(1+dL)}{L},\frac{d\rho}{1+\frac{dL}{2(1+dL)}},\frac{\rho^{2}+d^{2}}{2n}\big\}$, then $\psi(\delta)>0$ and
+$\frac{D_{\operatorname{tail}}}{D_{\operatorname{loc}}}\leq\left[\frac{\mathcal{K}_{\operatorname{geom}}}{\mathcal{K}_{\operatorname{err}}(\delta)}\right]\left(\frac{\delta}{\psi(\delta)}\right)\delta^{-\frac{(n-1)}{2}}e^{-\frac{\rho^{2}}{2\delta}},$
+where $\mathcal{K}_{\operatorname{err}}(\delta)$ and $\psi(\delta)$ are given in Claim 2. Moreover,
+$\mathcal{K}_{\operatorname{geom}}$ $=\frac{4\pi^{n/2}}{\Gamma\left(\frac{n}{2}\right)C_{\operatorname{vol}}(n)}(\rho^{2}+d^{2})^{\frac{n-2}{2}}(1+\rho L)^{\frac{n-1}{2}}.$
+Proof of Claim 3: We observe that for $0<\delta\leq\delta_{0}$,
+$\omega(z)$ $=\frac{\delta A(z)}{A(z)^{2}+\delta}-\frac{\delta}{\rho+d}\exp\left(-\frac{\left[(\rho+d)^{2}-A(z)^{2}\right]}{2\delta}\right)$
+$\geq\delta\left(\frac{A_{\operatorname{max}}}{A_{\operatorname{max}}^{2}+\delta}-\frac{1}{e(\rho+d)}\right)>0,$
+which implies that $\psi(\delta)=\inf_{z\in\mathbb{B}^{n-1}_{r_{\delta}}}\omega(z)>0$. Finally, the result follows from Claims 1 and 2. ∎
+Claim 4: Let $\delta_{1}:=\min\left\{\delta_{0},\frac{d^{2}}{3},\frac{\rho^{2}}{2\ln\!\left(4\left(1+\frac{\rho}{d}\right)\right)}\right\}$. Then, for all $\delta\in(0,\delta_{1})$, one has
+$\psi(\delta)\geq\frac{\delta}{2d}.$
+Proof of Claim 4: From the definition of $\omega(z)=T_{1}(z)-T_{2}(z)$, where
+$T_{1}(z)=\frac{\delta A(z)}{A(z)^{2}+\delta}\quad\textrm{ and }\quad T_{2}(z)=\frac{\delta}{\rho+d}\exp\left(-\frac{\left[(\rho+d)^{2}-A(z)^{2}\right]}{2\delta}\right)$
+Hence, on the one hand, if $\delta\leq d^{2}/3$, then $T_{1}(z)\geq\frac{3\delta}{4d}$.
+On the other hand, if $\delta\leq\frac{\rho^{2}}{2\ln\left(4(1+\frac{\rho}{d})\right)}$, then $T_{2}(z)\leq\frac{\delta}{4d}$.
+Finally,
+$\psi(\delta)=\inf_{z\in\mathbb{B}^{n-1}_{r_{\delta}}}(T_{1}(z)-T_{2}(z))\geq\frac{3\delta}{4d}-\frac{\delta}{4d}=\frac{\delta}{2d}.$
+With respect to the tangential/normal splitting induced by $y=(z,t)$, we write $N_{\operatorname{loc}}=(N_{\operatorname{loc}}^{\tau},N_{\operatorname{loc}}^{\nu})$ and $N_{\operatorname{tail}}=(N_{\operatorname{tail}}^{\tau},N_{\operatorname{tail}}^{\nu})$.
+Hence,
+$N_{\operatorname{loc}}^{\tau}$ $=\int_{\mathbb{B}_{\rho}^{n-1}}ze^{-\frac{\|z\|^{2}}{2\delta}}\left(\int_{-\rho}^{h(z)}e^{-\frac{(t-d)^{2}}{2\delta}}\,dt\right)\,dz$
+$N_{\operatorname{loc}}^{\nu}$ $=\int_{\mathbb{B}_{\rho}^{n-1}}e^{-\frac{\|z\|^{2}}{2\delta}}\underbrace{\left(\int_{-\rho}^{h(z)}t\,e^{-\frac{(t-d)^{2}}{2\delta}}\,dt\right)}_{=:I_{\nu}(z)}\,dz$
+Moreover, we write $p_{\delta}(x)=(p_{\delta}^{\tau}(x),p_{\delta}^{\nu}(x))$, where
+(13) $p_{\delta}^{\tau}(x)=\frac{N_{\operatorname{loc}}^{\tau}+N_{\operatorname{tail}}^{\tau}}{D_{\operatorname{loc}}+D_{\operatorname{tail}}}\quad p_{\delta}^{\nu}(x)=\frac{N_{\operatorname{loc}}^{\nu}+N_{\operatorname{tail}}^{\nu}}{D_{\operatorname{loc}}+D_{\operatorname{tail}}}.$
+Step 3: Bounding $N_{\operatorname{tail}}$.
+Claim 5:
+For all $0<\delta\leq\frac{d^{2}+\rho^{2}}{2n}$, one has
+$\|N_{\operatorname{tail}}\|\leq\frac{4\pi^{n/2}}{\Gamma\left(\frac{n}{2}\right)}\sqrt{\rho^{2}+d^{2}}(\rho^{2}+d^{2})^{\frac{n-2}{2}}\delta e^{-\frac{\rho^{2}+d^{2}}{2\delta}}.$
+Proof of Claim 5: From (12), one has
+$\|N_{\operatorname{tail}}\|$ $=\left\|\int_{C_{\operatorname{tail}}}ye^{-\frac{\|y-x\|^{2}}{2\delta}}\,dy\right\|$
+$\leq\int_{C_{\operatorname{tail}}}\|y\|e^{-\frac{\|y-x\|^{2}}{2\delta}}\,dy$
+$\leq\int_{\{u\in\mathbb{R}^{n}:\|u\|\geq\sqrt{\rho^{2}+d^{2}}\}}(\|u\|+d)e^{-\frac{\|u\|^{2}}{2\delta}}du,$
+which, by virtue of Lemma A.2, implies the claim.
+Step 4: Bounding the normal component $N_{\operatorname{loc}}^{\nu}$ of $N_{\operatorname{loc}}$.
+Claim 6:
+The following formula holds:
+(14) $N_{\operatorname{loc}}^{\nu}=d\,D_{\operatorname{loc}}-\delta I_{\operatorname{num}}+\delta\,e^{-\frac{(\rho+d)^{2}}{2\delta}}\int_{\mathbb{B}_{\rho}^{n-1}}e^{-\frac{\|z\|^{2}}{2\delta}}\,dz,$
+where $I_{\operatorname{num}}:=\int_{\mathbb{B}_{\rho}^{n-1}}e^{-\frac{\|z\|^{2}+A(z)^{2}}{2\delta}}\,dz$.
+Proof of Claim 6: Recall that $N_{\operatorname{loc}}^{\nu}=\int_{\mathbb{B}_{\rho}^{n-1}}e^{-\frac{\|z\|^{2}}{2\delta}}I_{\nu}(z)\,dz$ with
+$I_{\nu}(z)=\int_{-\rho}^{h(z)}t\,e^{-\frac{(t-d)^{2}}{2\delta}}\,dt=\underbrace{\int_{-(\rho+d)}^{-A(z)}ue^{-\frac{u^{2}}{2\delta}}\,du}_{\operatorname{Term}_{1}(z)}+d\underbrace{\int_{-(\rho+d)}^{-A(z)}e^{-\frac{u^{2}}{2\delta}}\,du}_{\operatorname{Term}_{2}(z)}.$
+Moreover, $\operatorname{Term}_{1}(z)=-\delta\big(e^{-\frac{A(z)^{2}}{2\delta}}-e^{-\frac{(\rho+d)^{2}}{2\delta}}\big)$ and $\operatorname{Term}_{2}(z)=J(z)$. Then,
+$I_{\nu}(z)=-\delta e^{-\frac{A(z)^{2}}{2\delta}}+\delta e^{-\frac{(\rho+d)^{2}}{2\delta}}+dJ(z),$
+which implies that
+$N_{\operatorname{loc}}^{\nu}$ $=\int_{\mathbb{B}_{\rho}^{n-1}}e^{-\frac{\|z\|^{2}}{2\delta}}\left[d\,J(z)-\delta e^{-\frac{A(z)^{2}}{2\delta}}+\delta e^{-\frac{(\rho+d)^{2}}{2\delta}}\right]\,dz$
+$=d\,D_{\operatorname{loc}}-\delta I_{\operatorname{num}}+\delta\,e^{-\frac{(\rho+d)^{2}}{2\delta}}\int_{\mathbb{B}_{\rho}^{n-1}}e^{-\frac{\|z\|^{2}}{2\delta}}\,dz,$
+which proves the claim. ∎
+We observe that the third term in the right-hand side of equality (14) is $\mathcal{O}(e^{-c/\delta})$, $D_{\operatorname{tail}}=\mathcal{O}(\delta e^{-c/\delta})$ (see Claim 1), and $\|N_{\operatorname{tail}}\|=\mathcal{O}(\delta e^{-c/\delta})$ (see Claim 5).
+Hence, informally speaking, by Claim 6, we get
+$p_{\delta}^{\nu}(x)=\frac{N_{\operatorname{loc}}^{\nu}+N_{\operatorname{tail}}^{\nu}}{D_{\operatorname{loc}}+D_{\operatorname{tail}}}\approx\frac{N_{\operatorname{loc}}^{\nu}}{D_{\operatorname{loc}}}\approx d-\delta\frac{I_{\operatorname{num}}}{D_{\operatorname{loc}}}.$
+Therefore, we consider the estimator
+$\tilde{p}_{\delta}^{\nu}:=d-\delta\frac{I_{\operatorname{num}}}{D_{\operatorname{loc}}}.$
+Let us consider the measure $\eta_{\delta}$ defined on $\mathbb{B}^{n-1}_{\rho}$ whose density is proportional to
+$e^{-\frac{\|z\|^{2}+(d-h(z))^{2}}{2\delta}}.$
+Claim 7: Assume that $0<\delta\leq\min\left\{\rho^{2}(1+dL),\frac{2\rho(1+dL)}{L},\frac{d\rho}{1+\frac{dL}{2(1+dL)}}\right\}$.
+Then there exists a function $r(A(z),\delta)$ such that $0<r(A(z),\delta)<\frac{\delta}{A(z)^{2}}$ and
+$\frac{I_{\operatorname{num}}}{D_{\operatorname{loc}}}=\frac{1}{\delta\mathbb{E}_{\eta_{\delta}}\left[\frac{1-r(A(z),\delta)}{A(z)}\right]-\Delta(\delta)}\quad\textrm{ for all }0<\delta\leq\delta_{1},$
+where
+$0\leq\Delta(\delta)\leq\frac{(1+dL)^{\frac{n-1}{2}}}{\mathcal{K}_{\operatorname{err}}\cdot(\rho+d)}\rho^{n-1}\frac{\operatorname{Vol}(\mathbb{B}^{n-1})}{\int_{\mathbb{B}^{n-1}}e^{-\frac{\|v\|^{2}}{2}}dv}\cdot\delta^{\frac{3-n}{2}}e^{-\frac{\rho^{2}+2d\rho}{2\delta}}=\mathcal{O}(\delta^{\frac{3-n}{2}}e^{-c/\delta}).$
+Proof of Claim 7: Let us observe that
+$J(z)$ $=\int_{A(z)}^{d+\rho}e^{-\frac{t^{2}}{2\delta}}\,dt$
+$=\int_{A(z)}^{\infty}e^{-\frac{t^{2}}{2\delta}}\,dt-\int_{d+\rho}^{\infty}e^{-\frac{t^{2}}{2\delta}}\,dt$
+$=\frac{\delta}{A(z)}e^{-\frac{A(z)^{2}}{2\delta}}\big(1-\underbrace{A(z)e^{\frac{A(z)^{2}}{2\delta}}\int_{A(z)}^{\infty}\frac{1}{t^{2}}e^{-\frac{t^{2}}{2\delta}}\,dt}_{=:r(A(z),\delta)}\big)-\int_{d+\rho}^{\infty}e^{-\frac{t^{2}}{2\delta}}\,dt.$
+Moreover, since $t\geq A(z)$ on $A(z),\infty)$, we have $1/t^{2}\leq 1/A(z)^{2}$. Using this bound and the standard Gaussian tail estimate (obtained by integration by parts), for all $w>0$
+(15) $\int_{w}^{\infty}e^{-\frac{t^{2}}{2\delta}}\,dt\leq\frac{\delta}{w}e^{-\frac{w^{2}}{2\delta}},$
+we obtain
+$0<r(A(z),\delta)$ $=A(z)e^{\frac{A(z)^{2}}{2\delta}}\int_{A(z)}^{\infty}\frac{1}{t^{2}}e^{-\frac{t^{2}}{2\delta}}\,dt$
+$\leq A(z)e^{\frac{A(z)^{2}}{2\delta}}\cdot\frac{1}{A(z)^{2}}\int_{A(z)}^{\infty}e^{-\frac{t^{2}}{2\delta}}\,dt$
+$\leq A(z)e^{\frac{A(z)^{2}}{2\delta}}\cdot\frac{1}{A(z)^{2}}\left(\frac{\delta}{A(z)}e^{-\frac{A(z)^{2}}{2\delta}}\right)$
+$=\frac{\delta}{A(z)^{2}}.$
+Therefore,
+$\frac{I_{\operatorname{num}}}{D_{\operatorname{loc}}}$ $=\frac{\int_{\mathbb{B}^{n-1}_{\rho}}e^{-\frac{\|z\|^{2}+A(z)^{2}}{2\delta}}\,dz}{\int_{\mathbb{B}^{n-1}_{\rho}}e^{-\frac{\|z\|^{2}}{2\delta}}J(z)\,dz}$
+$=\frac{1}{\delta\mathbb{E}_{\eta_{\delta}}\left[\frac{1-r(A(z),\delta)}{A(z)}\right]-\Delta(\delta)},$
+where
+$\Delta(\delta):=\frac{\int_{\rho+d}^{\infty}e^{-\frac{u^{2}}{2\delta}}\,du\cdot\int_{\mathbb{B}^{n-1}_{\rho}}e^{-\frac{\|z\|^{2}}{2\delta}}\,dz}{\int_{\mathbb{B}^{n-1}_{\rho}}e^{-\frac{\|z\|^{2}+A(z)^{2}}{2\delta}}\,dz}.$
+On the one hand, by inequality (15), we observe that the numerator of $\Delta(\delta)$ is bounded by
+$\frac{\delta}{\rho+d}e^{-\frac{(\rho+d)^{2}}{2\delta}}\operatorname{Vol}(\mathbb{B}^{n-1}_{\rho}).$
+On the other hand, the analysis for the denominator is similar to the analysis of $D_{\operatorname{loc}}$ (see Claim 2). Hence,
+$\int_{\mathbb{B}^{n-1}_{\rho}}e^{-\frac{\|z\|^{2}+A(z)^{2}}{2\delta}}\,dz\geq\mathcal{K}_{\operatorname{err}}\cdot e^{-\frac{d^{2}}{2\delta}}\left(\frac{\delta}{1+dL}\right)^{\frac{n-1}{2}}\int_{\mathbb{B}^{n-1}}e^{-\frac{\|v\|^{2}}{2}}dv.$
+Combining the above inequalities, we get the claim. ∎
+Claim 8: We have
+$\tilde{p}_{\delta}^{\nu}=\mathbb{E}_{\eta_{\delta}}[h(z)]-d\,\mathbb{E}_{\eta_{\delta}}\!\left[r(A(z),\delta)\right]+O(\delta^{2})\textrm{ as }\delta\to 0^{+}.$
+Moreover,
+$|\tilde{p}_{\delta}^{\nu}|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{2})\textrm{ as }\delta\to 0^{+}.$
+Proof of Claim 8: Let $g_{\delta}(z)=\frac{1-r(A(z),\delta)}{A(z)}$. Then, by virtue of Claim 7, we get
+$\tilde{p}_{\delta}^{\nu}$ $=d-\delta\frac{I_{\operatorname{num}}}{D_{\operatorname{loc}}}=d-\frac{\delta}{\delta\mathbb{E}_{\eta_{\delta}}[g_{\delta}(z)]}+O(\delta^{\frac{1-n}{2}}e^{-c/\delta}).$
+Moreover, by Taylor’s theorem,
+$\frac{1}{\mathbb{E}_{\eta_{\delta}}[g_{\delta}(z)]}=\frac{1}{\frac{1}{d}+\mathbb{E}_{\eta_{\delta}}[g_{\delta}(z)-\frac{1}{d}]}=d-d^{2}\mathbb{E}_{\eta_{\delta}}[g_{\delta}(z)-\frac{1}{d}]+O(\delta^{2}).$
+Therefore,
+$\tilde{p}_{\delta}^{\nu}=\mathbb{E}_{\eta_{\delta}}[h(z)]-d\,\mathbb{E}_{\eta_{\delta}}[r(A(z),\delta)]+O(\delta^{2}).$
+Finally, using that $|h(z)|\leq\frac{L}{2}\|z\|^{2}$ and $\mathbb{E}_{\eta_{\delta}}[\|z\|^{2}]=(n-1)\delta+O(\delta^{2})$, we obtain that
+$|\tilde{p}_{\delta}^{\nu}|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{2}),$
+which proves the claim.
+Step 4: Bounding the tangential component $N_{\operatorname{loc}}^{\tau}$ of $N_{\operatorname{loc}}$.
+Claim 9:
+For all $0<\delta\leq\delta_{1}$
+$\|N_{\operatorname{loc}}^{\tau}\|\leq\frac{M(n^{2}-1)}{6}(2\pi\delta)^{\frac{n-1}{2}}\delta^{2}e^{-\frac{d^{2}}{2\delta}}.$
+Proof of Claim 9: Recall that
+$N_{\operatorname{loc}}^{\tau}=\int_{\mathbb{B}_{\rho}^{n-1}}ze^{-\frac{\|z\|^{2}}{2\delta}}J(z)dz.$
+Since $\mathbb{B}_{\rho}^{n-1}$ is symmetric and $e^{-\frac{\|z\|^{2}}{2\delta}}$ is even,
+$N_{\operatorname{loc}}^{\tau}=\frac{1}{2}\int_{\mathbb{B}_{\rho}^{n-1}}ze^{-\frac{\|z\|^{2}}{2\delta}}(J(z)-J(-z))dz.$
+Hence
+$\|N_{\operatorname{loc}}^{\tau}\|\leq\frac{1}{2}\int_{\mathbb{B}_{\rho}^{n-1}}\|z\|e^{-\frac{\|z\|^{2}}{2\delta}}|J(z)-J(-z)|dz.$
+Set $F(u):=\int_{-\rho}^{u}e^{-\frac{(t-d)^{2}}{2\delta}}dt$.
+Hence, $J(z)=F(h(z))$ and $F^{\prime}(u)=e^{-\frac{(u-d)^{2}}{2\delta}}$.
+By the mean value theorem, for each $z$ there exists $\theta_{z}$ between $h(z)$ and h(z)h(-z) such that
+$J(z)-J(-z)=F^{\prime}(\theta_{z})(h(z)-h(-z)).$
+Moreover, since $h(z)\leq 0$ and $h(-z)\leq 0$, we have $\theta_{z}\leq 0$, hence $(\theta_{z}-d)^{2}\geq d^{2}$ and thus
+$|F^{\prime}(\theta_{z})|\leq e^{-\frac{d^{2}}{2\delta}}.$
+Therefore,
+$|J(z)-J(-z)|\leq e^{-\frac{d^{2}}{2\delta}}|h(z)-h(-z)|.$
+From (10), we get that
+$h(z)=\frac{1}{2}\langle Hz,z\rangle+r(z),\quad|r(z)|\leq\frac{M}{6}\|z\|^{3}.$
+Since the quadratic term is even,
+h(z)h(z)=$r(z)$,h(z)-h(-z)=r(z)-r(-z),
+and hence $|J(z)-J(-z)|\leq\frac{M}{3}\|z\|^{3}e^{-\frac{d^{2}}{2\delta}}$. Using the above estimates, we get
+$\|N_{\operatorname{loc}}^{\tau}\|\leq\frac{M}{6}e^{-\frac{d^{2}}{2\delta}}\int_{\mathbb{B}_{\rho}^{n-1}}\|z\|^{4}e^{-\frac{\|z\|^{2}}{2\delta}}dz\leq\frac{M}{6}e^{-\frac{d^{2}}{2\delta}}\int_{\mathbb{R}^{n-1}}\|z\|^{4}e^{-\frac{\|z\|^{2}}{2\delta}}dz$
+If $W\sim\mathcal{N}(0,\delta I_{n-1})$, then
+$\int_{\mathbb{R}^{n-1}}\|z\|^{4}e^{-\frac{\|z\|^{2}}{2\delta}}dz=(2\pi\delta)^{\frac{n-1}{2}}\mathbb{E}\|W\|^{4},\quad\mathbb{E}\|W\|^{4}=\delta^{2}(n^{2}-1),$
+which yields the claim. ∎
+Claim 10: Under the assumptions of Theorem 4.3, we have
+$\|p_{\delta}^{\tau}(x)\|\leq\frac{M(n^{2}-1)d}{6}\delta+O(\delta^{2})\textrm{ as }\delta\to 0^{+}.$
+Proof of Claim 10: By (13),
+$\|p_{\delta}^{\tau}(x)\|\leq\frac{\|N_{\operatorname{loc}}^{\tau}\|}{D_{\operatorname{loc}}+D_{\operatorname{tail}}}+\frac{\|N_{\operatorname{tail}}\|}{D_{\operatorname{loc}}+D_{\operatorname{tail}}}\leq\frac{\|N_{\operatorname{loc}}^{\tau}\|}{D_{\operatorname{loc}}}+\frac{\|N_{\operatorname{tail}}\|}{D_{\operatorname{loc}}}.$
+Using Claim 9 and Claim 2-4, for $0<\delta\leq\delta_{1}$ we have
+$\frac{\|N_{\operatorname{loc}}^{\tau}\|}{D_{\operatorname{loc}}}$ $\leq\frac{\frac{M(n^{2}-1)}{6}(2\pi\delta)^{\frac{n-1}{2}}\delta^{2}e^{-\frac{d^{2}}{2\delta}}}{\psi(\delta)\mathcal{K}_{\rm err}(\delta)\frac{C_{\rm vol}(n)}{(1+dL)^{\frac{n-1}{2}}}\delta^{\frac{n-1}{2}}e^{-\frac{d^{2}}{2\delta}}}$
+$=\frac{M(n^{2}-1)}{6}\,\frac{(2\pi)^{\frac{n-1}{2}}(1+dL)^{\frac{n-1}{2}}}{\mathcal{K}_{\rm err}(\delta)C_{\rm vol}(n)}\cdot\frac{\delta^{2}}{\psi(\delta)}.$
+By Claim 4, $\psi(\delta)\geq\delta/(2d)$, hence
+$\frac{\|N_{\rm loc}^{\tau}\|}{D_{\rm loc}}\leq\frac{M(n^{2}-1)d}{3}\,\frac{(2\pi)^{\frac{n-1}{2}}(1+dL)^{\frac{n-1}{2}}}{\mathcal{K}_{\rm err}(\delta)C_{\rm vol}(n)}\cdot\delta.$
+Finally, Claim 5 and Claim 2–4 imply $\|N_{\rm tail}\|/D_{\rm loc}=O(e^{-c/\delta})$, which can be absorbed into $O(\delta^{2})$ as $\delta\to 0^{+}$. This completes the proof. ∎
+Step 5: Conclusion.
+Since $\operatorname{proj}_{C}(x)=0$ in the chosen coordinates, we have
+$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|=\|p_{\delta}(x)\|\leq|p_{\delta}^{\nu}(x)|+\|p_{\delta}^{\tau}(x)\|.$
+By Claim 8 (together with Claims 1,5,6 to control the tail and the exponentially small terms), we obtain
+$|p_{\delta}^{\nu}(x)|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{2}).$
+By Claim 10,
+$\|p_{\delta}^{\tau}(x)\|\leq C_{\tau}\,\delta+O(\delta^{2}),\qquad C_{\tau}=\frac{M(n^{2}-1)d}{3}\,\frac{(2\pi)^{\frac{n-1}{2}}(1+dL)^{\frac{n-1}{2}}}{C_{\rm vol}(n)}.$
+Combining the two estimates yields the theorem. ∎

### v3

@@
 1. Introduction
-Proximal mappings and metric projections are fundamental primitives in variational analysis and optimization.
+Proximal mappings and metric projections are fundamental concepts in variational analysis and optimization.
+Introduced by J.-J.
+Moreau in [9], proximal mappings have since become central tools in modern optimization, notably through proximal algorithms and operator-splitting methods and their many variants (see, e.g., [2]).
 Given a proper function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ and $\lambda>0$, the proximal point $\operatorname{prox}_{\lambda f}(x)$ is the unique minimizer of $y\mapsto f(y)+\frac{1}{2\lambda}\|x-y\|^{2}$ whenever $f$ is $\rho$-weakly convex and $\lambda\in(0,1/\rho)$.
@@
 This paper studies a family of stochastic (zeroth-order) estimators for proximal points and projections based on Gaussian perturbations and exponential reweighting.
-For $\delta>0$, define the barycenter
+For $\delta>0$ and $\lambda>0$, we define the barycenter
 (1) $m_{\delta}(x):=\frac{\mathbb{E}_{Y\sim\mathcal{N}(x,\delta\lambda I)}\!\big[Y\,\exp(-f(Y)/\delta)\big]}{\mathbb{E}_{Y\sim\mathcal{N}(x,\delta\lambda I)}\!\big[\exp(-f(Y)/\delta)\big]}.$
@@
 If $f\in C^{2}(\mathbb{R}^{n})$ is $\rho$-weakly convex and $\nabla^{2}f$ is globally Lipschitz with constant $L$, we establish the linear-rate refinement
-$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{3nL}{2\mu^{2}}\,\delta,\qquad\mu=\frac{1}{\lambda}-\rho,$
+$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{nL}{\mu^{2}}\,\delta,\qquad\mu=\frac{1}{\lambda}-\rho,$
 again for every $\delta>0$. This result quantifies a phenomenon suggested by the smoothing interpretations in [10, 11]: additional curvature regularity improves the bias of the barycentric estimator.
@@
 The closed unit ball in $\mathbb{R}^{n}$ is denoted by $\mathbb{B}$.
+The open ball of radius $R$ centered at $x$ is denoted by $B_{R}(x)$.
+For simplicity, the open ball of radius $R$ centered at 0 is denoted by $B_{R}$.
+Moreover, $\mathbb{S}^{n-1}$ denotes the unit sphere in $\mathbb{R}^{n}$.
 We use the convention $\exp(+\infty)=+\infty$ and $\exp(-\infty)=0$.
-For a ser $A\subset\mathbb{R}^{n}$, the indicator function of $A$ is defined by
+For a set $A\subset\mathbb{R}^{n}$, the indicator function of $A$ is defined by
 $\iota_{A}(x):=\left\{\begin{array}[]{cc}0&\text{ if }x\in A,\\ +\infty&\text{ if }x\notin A.\end{array}\right$.
@@
 Moreover, if $A$ is nonempty, closed and convex, the projection of $x\in\mathbb{R}^{n}$ onto $A$ is denoted by $\operatorname{proj}_{A}(x)$.
-Given a function $g\colon\mathbb{R}^{n}\rightarrow\mathbb{\mathbb{R}}\cup\{+\infty\}$, the domain of $g$ is
+Given a function $g\colon\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}$, the domain of $g$ is
 $\operatorname{dom}(g):=\{x\in\mathbb{R}^{n}:g(x)<+\infty\}.$
@@
 For $\mu>0$, we say that $g$ is $\mu$-strongly convex if $g-\frac{\mu}{2}\|\cdot\|^{2}$ is convex.
-The Moreau envelope of index $\lambda>0$ of a function $g\colon\mathbb{R}^{n}\to{\mathbb{R}}_{\infty}$ is denoted by $\operatorname{e}_{\lambda}g\colon\mathbb{R}^{n}\to\mathbb{R}$ and is defined by
+The Moreau envelope of index $\lambda>0$ of a function $g\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is denoted by $\operatorname{e}_{\lambda}g\colon\mathbb{R}^{n}\to\mathbb{R}$ and is defined by
 $\operatorname{e}_{\lambda}g(x):=\inf_{y\in\mathbb{R}^{n}}\left(g(y)+\frac{1}{2\lambda}\|x-y\|^{2}\right)\textrm{ for all }x\in\mathbb{R}^{n}.$
@@
 $\mathbb{E}_{\sigma_{\delta}}[w(y)]=\frac{\int_{\mathbb{R}^{n}}w(y)e^{-g(y)/\delta}dy}{\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy}$
-the expectation of a given function $w$ with respect to this probability measure.
-The following result is known as Fradelizi’s inequality (see [6])
-Lemma 2.1.
-Let $h:\mathbb{R}^{n}\to[0,\infty)$ be a log-concave function, not identically zero, such that $\int_{\mathbb{R}^{n}}h(y)\,dy<\infty$. Define its barycenter by
-$b_{h}:=\frac{\int_{\mathbb{R}^{n}}y\,h(y)\,dy}{\int_{\mathbb{R}^{n}}h(y)\,dy}.$
-Then $b_{h}$ is well-defined and
-$h(b_{h})\ \geq\ e^{-n}\,\sup_{z\in\mathbb{R}^{n}}h(z).$
-3. Nonsmooth results
-Let $f\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ be a $\rho$-weakly convex function with $\rho\geq 0$. The aim of this section is to prove that, for any $0<\lambda<1/\rho$,
-$m_{\delta}(x):=\frac{\mathbb{E}_{y\sim\mathcal{N}(x,\delta\lambda I)}[y\cdot\exp(-f(y)/\delta)]}{\mathbb{E}_{y\sim\mathcal{N}(x,\delta\lambda I)}[\exp(-f(y)/\delta)]}\to\operatorname{prox}_{\lambda f}(x)\textrm{ as }\delta\to 0^{+}.$
-The next result provides a rate of convergence of the barycenter toward the projection.
+the expectation of a given function $w$ with respect to the probability measure $\sigma_{\delta}$, which is formally defined for every Lebesgue-measurable set $A$ as
+$\sigma_{\delta}(A):=\frac{\int_{A}e^{-g(y)/\delta}dy}{\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy}.$
+3. Rates of Convergence for Nonsmooth Functions
+Let $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ be a $\rho$-weakly convex function with $\rho\geq 0$. The goal of this section is to establish an explicit convergence rate for the map
+$m_{\delta}(x):=\frac{\mathbb{E}_{y\sim\mathcal{N}(x,\delta\lambda I)}\!\bigl[y\,\exp(-f(y)/\delta)\bigr]}{\mathbb{E}_{y\sim\mathcal{N}(x,\delta\lambda I)}\!\bigl[\exp(-f(y)/\delta)\bigr]},$
+and to quantify its approximation of the proximal point:
+$m_{\delta}(x)\;\longrightarrow\;\operatorname{prox}_{\lambda f}(x)\qquad\text{as }\delta\downarrow 0.$
+It is straightforward to verify that if $f$ is $\rho$-weakly convex for some $\rho\geq 0$ and $\operatorname{dom}f$ has nonempty interior, then $m_{\delta}$ is well-defined for every $\delta>0$ and $0<\lambda<1/\rho$.
+The following result provides a convergence rate for this barycenter toward the proximal point.
 Theorem 3.1.
@@
 $F(y):=(y-\operatorname{prox}_{\lambda f}(x))e^{-g(y)/\delta}.$
-Claim 1: $\mathbb{E}_{\sigma_{\delta}}\left[\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle\right]\leq n\delta$.
-Proof of Claim 1: Since $y\mapsto g(y)$ is proper, lsc, and convex, its domain is convex; in particular, $D:=\operatorname{dom}f=\operatorname{dom}g$ is convex with nonempty interior (by assumption).
+Now, let us show that $\mathbb{E}_{\sigma_{\delta}}\left[\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle\right]\leq n\delta$.
+Indeed, since $y\mapsto g(y)$ is proper, lsc, and convex, its domain is convex; in particular, $D:=\operatorname{dom}f=\operatorname{dom}g$ is convex with nonempty interior (by assumption).
 On $\operatorname{int}(D)$, every proper lsc convex function is locally Lipschitz, hence differentiable a.e.
@@
 $\operatorname{div}F(y)=ne^{-g(y)/\delta}-\frac{1}{\delta}\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle e^{-g(y)/\delta}.$
-Fix $R>0$ and set $\Omega_{R}:=\operatorname{int}(D)\cap\mathbb{B}_{R}$. Since
-$\|y-\operatorname{prox}_{\lambda f}(x)\|\leq R+\|\operatorname{prox}_{\lambda f}(x)\|=:C_{R}\textrm{ on }\mathbb{B}_{R}$
-and $e^{-g(y)/\delta}\leq e^{-\inf_{\mathbb{B}_{R}}g/\delta}=:M_{R,\delta}<\infty$, it follows that for a.e. $y\in\Omega_{R}$,
+Fix $R>0$ and set $\Omega_{R}:=\operatorname{int}(D)\cap B_{R}$. Since
+$\|y-\operatorname{prox}_{\lambda f}(x)\|\leq R+\|\operatorname{prox}_{\lambda f}(x)\|=:C_{R}\textrm{ on }B_{R}$
+and $e^{-g(y)/\delta}\leq e^{-\inf_{B_{R}}g/\delta}=:M_{R,\delta}<\infty$, it follows that for a.e. $y\in\Omega_{R}$,
 $|\operatorname{div}F(y)|\leq nM_{R,\delta}+\frac{M_{R,\delta}C_{R}}{\delta}\|\nabla g(y)\|,$
 Since $\nabla g\in L^{1}_{\operatorname{loc}}(\operatorname{int}(D);\mathbb{R}^{n})$, we have $\nabla g\in L^{1}(\Omega_{R};\mathbb{R}^{n})$, and then $\operatorname{div}F\in L^{1}(\Omega_{R})$.
-For $R>0$, apply the divergence theorem on $\Omega_{R}$:
+For $R>0$, apply the divergence theorem to $\Omega_{R}$:
 $\int_{\Omega_{R}}\operatorname{div}F(y)dy$ $=\int_{\partial\Omega_{R}}F\cdot\nu\,d\mathcal{H}^{n-1}.$
 Moreover, by Lemma A.1, we can select $R_{k}\to\infty$ such that
-$\int_{\partial\Omega_{R_{k}}}\langle F,\nu\rangle\,d\mathcal{H}^{n-1}=\underbrace{\int_{\partial D\cap\mathbb{B}_{R_{k}}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}}_{=:I_{1}}+\underbrace{\int_{\partial\mathbb{B}_{R_{k}}\cap D}F\cdot\nu_{\mathbb{B}_{R_{k}}}\,d\mathcal{H}^{n-1}}_{=:I_{2}}.$
+$\int_{\partial\Omega_{R_{k}}}F\cdot\nu\,d\mathcal{H}^{n-1}=\underbrace{\int_{\partial D\cap{B}_{R_{k}}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}}_{=:I_{1}}+\underbrace{\int_{\partial{B}_{R_{k}}\cap D}F\cdot\nu_{{B}_{R_{k}}}\,d\mathcal{H}^{n-1}}_{=:I_{2}}.$
 Now, on the one hand, since $D$ is convex and $\operatorname{prox}_{\lambda f}(x)\in D$, for $\mathcal{H}^{n-1}$-a.e. $y\in\partial D$ one has
-$\langle y-\operatorname{prox}_{\lambda f}(x),\nu_{D}(y)\rangle\geq 0\quad\Rightarrow\quad\langle F(y),\nu_{D}(y)\rangle\geq 0.$
-Hence the $\partial D$-term $I_{1}$ is nonnegative. On the other hand, using (2), the term $I_{1}$ is bounded in absolutely value by
+$\langle y-\operatorname{prox}_{\lambda f}(x),\nu_{D}(y)\rangle\geq 0\quad\Rightarrow\quad F(y)\cdot\nu_{D}(y)\geq 0.$
+Hence the $\partial D$-term $I_{1}$ is nonnegative. On the other hand, using (2), the term $I_{2}$ is bounded in absolute value by
 $CR_{k}^{n}\exp\left(-\frac{\mu}{2\delta}(R_{k}-\|\operatorname{prox}_{\lambda f}(x)\|)^{2}\right)\to 0\textrm{ as }k\to\infty,$
-hence the $\partial\mathbb{B}_{R_{k}}$-term $I_{2}$ vanishes as $k\to\infty$. Therefore,
-$0$ $\leq\limsup_{k\to\infty}\int_{\mathbb{B}_{R_{k}}\cap\partial D}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}$
-$=\limsup_{k\to\infty}\Big(\int_{\partial D\cap\mathbb{B}_{R_{k}}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\partial\mathbb{B}_{R_{k}}\cap D}F\cdot\nu_{\mathbb{B}_{R_{k}}}\,d\mathcal{H}^{n-1}\Big)$
-$=\limsup_{k\to\infty}\int_{\partial\Omega_{R_{k}}}F\,\nu\,d\mathcal{H}^{n-1}$
+hence the $\partial{B}_{R_{k}}$-term $I_{2}$ vanishes as $k\to\infty$. Therefore,
+$0$ $\leq\limsup_{k\to\infty}\int_{B_{R_{k}}\cap\partial D}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}$
+$=\limsup_{k\to\infty}\Big(\int_{\partial D\cap{B}_{R_{k}}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\partial{B}_{R_{k}}\cap D}F\cdot\nu_{{B}_{R_{k}}}\,d\mathcal{H}^{n-1}\Big)$
+$=\limsup_{k\to\infty}\int_{\partial\Omega_{R_{k}}}F\cdot\nu\,d\mathcal{H}^{n-1}$
 $=\limsup_{k\to\infty}\int_{\Omega_{R_{k}}}\operatorname{div}F(y)dy$
@@
 $=n\int_{D}e^{-g(y)/\delta}dy-\frac{1}{\delta}\int_{D}\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle e^{-g(y)/\delta}dy.$
-Finally, dividing by $\int_{D}e^{-g(y)/\delta}dy$ and rearranging, we obtain the claim.∎
-Next, by strong convexity and almost-everywhere differentiability, we obtain
+Finally, dividing by $\int_{D}e^{-g(y)/\delta}dy$ and rearranging, we obtain the estimate. Next, by strong convexity and almost-everywhere differentiability, we obtain
 $\langle\nabla g(y),y-\operatorname{prox}_{\lambda f}(x)\rangle\geq\mu\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}.$
@@
 $\mu\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\right]\leq\mathbb{E}_{\sigma_{\delta}}\left[\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle\right]=n\delta,$
-hence $\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\right]\leq n\delta/\mu$. Finally, by Jensen inequality,
+hence $\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\right]\leq n\delta/\mu$. Finally, by Jensen’s inequality,
 $\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2}$ $\leq\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|\right]^{2}$
@@
 This shows that the order of convergence in Theorem 3.1 is sharp.
-The following example shows that the bound obtained in Theorem 3.1 is sharp.
+The following example shows that the bound obtained in Theorem 3.1 is optimal.
 Example 3.2.
-Fix $n\geq 2$, take $\lambda>0$, set $\mu=1/\lambda$, and let $x=0$. For $\alpha\in(0,\pi/2)$, define the circular cone
-$K_{\alpha}:=\{y\in\mathbb{R}^{n}\colon\langle y,e_{1}\rangle\geq\|y\|\,\cos\alpha\}.$
-Let $f=\iota_{K_{\alpha}}$ be the indicator function of $K_{\alpha}$. Then $\operatorname{prox}_{\lambda f}(0)=0$. Moreover,
-$\|p_{\delta}(0)-\operatorname{prox}_{\lambda f}(0)\|=\|m_{\delta}(0)\|=\sqrt{\frac{\delta}{\mu}}\mathbb{E}[\chi_{n}]m_{\alpha,n},$
+Fix $n\geq 2$, take $\lambda>0$, set $\mu:=1/\lambda$, and let $x=0$. For $\alpha\in(0,\pi/2)$, define the circular cone
+$K_{\alpha}:=\{y\in\mathbb{R}^{n}\colon\langle y,e_{1}\rangle\geq\|y\|\,\cos\alpha\},$
+where $e_{1}$ denotes the first canonical basis vector. Let $f=\iota_{K_{\alpha}}$ be the indicator function of $K_{\alpha}$. Then $\operatorname{prox}_{\lambda f}(0)=0$. Moreover,
+$\|m_{\delta}(0)-\operatorname{prox}_{\lambda f}(0)\|=\|m_{\delta}(0)\|=\sqrt{\frac{\delta}{\mu}}\mathbb{E}[\chi_{n}]\mathfrak{m}_{\alpha,n},$
 where $\chi_{n}$ denotes the $\chi$-distribution with $n$ degrees of freedom and
-$m_{\alpha,n}:=\mathbb{E}\left[U_{1}\mid U_{1}\geq\cos\alpha\right],$
+$\mathfrak{m}_{\alpha,n}:=\mathbb{E}\left[U_{1}\mid U_{1}\geq\cos\alpha\right],$
 with $U$ uniformly distributed on $\mathbb{S}^{n-1}$ and $U_{1}=\langle U,e_{1}\rangle$.
-In particular, since $U_{1}\in[\cos\alpha,1]$ on the conditioning event, we have $m_{\alpha,n}\in[\cos\alpha,1]$ and hence $m_{\alpha,n}\to 1$ as $\alpha\searrow 0$.
+In particular, since $U_{1}\in[\cos\alpha,1]$ on the conditioning event, we have $\mathfrak{m}_{\alpha,n}\in[\cos\alpha,1]$ and hence $\mathfrak{m}_{\alpha,n}\to 1$ as $\alpha\searrow 0$.
 Furthermore,
-$\mathbb{E}[\chi_{n}]\sqrt{n}\to 1\textrm{ as }n\to\infty.$
+$\frac{\mathbb{E}[\chi_{n}]}{\sqrt{n}}\to 1\textrm{ as }n\to\infty.$
 Consequently, for every $\varepsilon>0$ there exist $n$ and $\alpha$ such that, for all $\delta>0$,
-$\|p_{\delta}(0)-\operatorname{prox}_{\lambda f}(0)\|\geq(1-\varepsilon)\sqrt{\frac{n\delta}{\mu}}.$
+$\|m_{\delta}(0)-\operatorname{prox}_{\lambda f}(0)\|\geq(1-\varepsilon)\sqrt{\frac{n\delta}{\mu}}.$
 This shows that the rate of convergence in Theorem 3.1 is sharp.
@@
 Fix $x\in\mathbb{R}^{n}$ and $0<\lambda<1/\rho$, and set $\mu:=\frac{1}{\lambda}-\rho>0$. Then, for every $\delta>0$,
-$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{3nL}{2\mu^{2}}\,\delta.$
+$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{nL}{\mu^{2}}\,\delta.$
 Proof.
@@
 $g(y):=f(y)+\frac{1}{2\lambda}\|x-y\|^{2}.$
-Since $f$ is $\rho$-weakly convex and $0<\lambda<1/\rho$, it follows that $g$ is $\mu$-strongly convex and
-(3) $g(y)\geq g(\operatorname{prox}_{\lambda f}(x))+\frac{\mu}{2}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\textrm{ for all }y\in\mathbb{R}^{n}.$
-From this inequality we get that $e^{-g/\delta}$ is integrable for every $\delta>0$, so $m_{\delta}(x)$ is well-defined.
-Given a function $w(\cdot)$, we define
-$\mathbb{E}_{\sigma_{\delta}}[w(y)]=\frac{\int_{\mathbb{R}^{n}}w(y)e^{-g(y)/\delta}dy}{\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy}$
+Since $f$ is $\rho$-weakly convex and $0<\lambda<1/\rho$, it follows that $g$ is $\mu$-strongly convex and that (2) holds.
+In particular, (2) implies that $e^{-g/\delta}$ is integrable for every $\delta>0$, so $m_{\delta}(x)$ is well-defined.
+Given a function $w(\cdot)$, recall that
+$\mathbb{E}_{\sigma_{\delta}}[w(y)]=\frac{\int_{\mathbb{R}^{n}}w(y)e^{-g(y)/\delta}dy}{\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy}.$
 Claim 1: $\mathbb{E}_{\sigma_{\delta}}[\nabla g(y)]=0$.
-Proof of Claim 1: We observe that
+Proof of Claim 1: First, note that the Lipschitz continuity of the Hessian of $f$ implies that $\nabla g$ grows at most quadratically. More precisely, there exist constants $a,b>0$ such that
+$\|\nabla g(y)\|\leq a\|y\|^{2}+b\qquad\text{for all }y\in\mathbb{R}^{n}.$
+In particular, $\nabla g$ is integrable with respect to the probability measure $\sigma_{\delta}$.
+Second, we observe that
 $\frac{\partial}{\partial y_{i}}\left(e^{-g(y)/\delta}\right)=-\frac{1}{\delta}\frac{\partial g}{\partial y_{i}}(y)e^{-g(y)/\delta}.$
 Moreover, from (2), it follows that
-$0\leq e^{-g(y)/\delta}\leq e^{-g(\operatorname{prox}_{\lambda f}(x))/\delta}\cdot e^{-\frac{\mu}{\delta}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}}\textrm{ for all }y\in\mathbb{R}^{n},$
+$0\leq e^{-g(y)/\delta}\leq e^{-g(\operatorname{prox}_{\lambda f}(x))/\delta}\cdot e^{-\frac{\mu}{2\delta}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}}\textrm{ for all }y\in\mathbb{R}^{n},$
 which implies that
@@
 Claim 4: For every $\delta>0$, we have
-$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{3nL}{2\mu^{2}}\delta.$
+$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{nL}{\mu^{2}}\delta.$
 Proof of Claim 4: On the one hand, from Claims 2 and 3, we have
@@
 $\leq\frac{L}{2\mu}\mathbb{E}_{\sigma_{\delta}}[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}],$
-where we have used that $H\succcurlyeq\mu I$ and $\|H^{-1}\|\leq 1/\mu$. Moreover,
+where we used that $H\succcurlyeq\mu I$, hence $\|H^{-1}\|_{2}\leq 1/\mu$, where $\|\cdot\|_{2}$ is the induced $\ell_{2}$-operator norm.
+Moreover, using that $\mathbb{E}_{\sigma_{\delta}}[y]=m_{\delta}(x)$, we get that
 (4) $\mathbb{E}_{\sigma_{\delta}}[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}]=\mathbb{E}_{\sigma_{\delta}}[\|y-m_{\delta}(x)\|^{2}]+\mathbb{E}_{\sigma_{\delta}}[\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2}].$
@@
 From Brascamp–Lieb inequality (see [3, Theorem 4.1]), we get that
-(6) $\operatorname{Cov}_{\sigma_{\delta}}(y)\preceq\frac{\delta}{\mu}I\quad\Rightarrow\quad\mathbb{E}_{\sigma_{\delta}}[\|y-m_{\delta}(x)\|^{2}]=\operatorname{tr}\left(\operatorname{Cov}_{\sigma_{\delta}}(y)\right)\leq\operatorname{tr}\left(\frac{\delta}{\mu}I\right)=\frac{n\delta}{\mu}.$
-Next, again since $h_{\delta}=e^{-g/\delta}$ is log-concave, applying the Lemma 2.1 to $h_{\delta}$ yields
-(7) $g(m_{\delta}(x))-g(\operatorname{prox}_{\lambda f}(x))\leq n\delta.$
-By $\mu$-strong convexity of $g$,
-$g(m_{\delta}(x))-g(\operatorname{prox}_{\lambda f}(x))\geq\frac{\mu}{2}\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2},$
-and combining with (7) gives
-$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2}\leq\frac{2n\delta}{\mu}.$
+(5) $\operatorname{Cov}_{\sigma_{\delta}}(y)\preceq\frac{\delta}{\mu}I\qquad\Rightarrow\qquad\mathbb{E}_{\sigma_{\delta}}[\|y-m_{\delta}(x)\|^{2}]=\operatorname{tr}\left(\operatorname{Cov}_{\sigma_{\delta}}(y)\right)\leq\operatorname{tr}\left(\frac{\delta}{\mu}I\right)=\frac{n\delta}{\mu},$
+where $\operatorname{Cov}_{\sigma_{\delta}}$ denotes the covariance matrix of a random vector under the probability measure $\sigma_{\delta}$. Moreover, Theorem 3.1 gives us the estimation
+$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2}\leq\frac{n\delta}{\mu}.$
 Plugging (5) and the above inequality into (4) yields
-$\mathbb{E}_{\sigma_{\delta}}[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}]\leq\frac{n\delta}{\mu}+\frac{2n\delta}{\mu}=\frac{3n\delta}{\mu}.$
+$\mathbb{E}_{\sigma_{\delta}}[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}]\leq\frac{n\delta}{\mu}+\frac{n\delta}{\mu}=\frac{2n\delta}{\mu}.$
 Finally, from the above inequality and (3), we get
-$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{L}{2\mu}\cdot\frac{3n\delta}{\mu}=\frac{3nL}{2\mu^{2}}\delta,$
+$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{L}{2\mu}\cdot\frac{2n\delta}{\mu}=\frac{nL}{\mu^{2}}\delta,$
 which proves Claim 4 and the theorem. ∎
@@
 Let $C\subset\mathbb{R}^{n}$ be a closed convex set and let $p\in\partial C$.
-We say that $\partial C$ admits a local $C^{2,1}$ chart at $p$ with constants $(\rho,L,M)$ if, after a rigid motion sending $p$ to 0 and the outward unit normal at $p$ to $e_{n}$, there exists $\rho>0$ and a concave function $h\in C^{2,1}(\mathbb{B}^{n-1}_{\rho})$ such that $h(0)=0$, $\nabla h(0)=0$, $\|\nabla^{2}h(z)\|\leq L$,
+We say that $\partial C$ admits a local $C^{2,1}$ chart at $p$ with constants $(\rho,L,M)$ (see Figure 1) if, after a rigid motion sending $p$ to 0 and the outward unit normal at $p$ to $e_{n}$, there exists $\rho>0$ and a concave function $h\in C^{2,1}(\mathbb{B}^{n-1}_{\rho})$ such that $h(0)=0$, $\nabla h(0)=0$, $\|\nabla^{2}h(z)\|\leq L$,
 $C\cap\big(\mathbb{B}^{n-1}_{\rho}\times(-\rho,\rho)\big)=\{(z,t)\in\mathbb{R}^{n-1}\times\mathbb{R}:\ t\leq h(z)\},$
@@
 Figure 1. Local $C^{2,1}$ chart of $\partial C$.
-In the case of a smooth convex set, we obtain the following improvement of Corollary 3.3, which provides the convergence rate
-$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|=O(\delta)\textrm{ as }\delta\to 0^{+}.$
+In the case of a smooth convex set, we obtain the following refinement of Corollary 3.3, which yields a better convergence rate of order $\delta$.
 Theorem 4.3.
@@
 Assume that $\partial C$ admits a local $C^{2,1}$ chart at $\operatorname{proj}_{C}(x)$ with constants $(\rho,L,M)$.
-Then there exist $\mathcal{C}>0$ and $\delta_{0}>0$ such for every $0<\delta<\delta_{0}$, we have
-$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|\leq$ $\,\mathcal{C}\delta+\mathcal{O}(\delta^{2}).$
+Then there exists $\delta_{0}>0$ such that for every $0<\delta<\delta_{0}$, we have
+$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|=$ $\mathcal{O}(\delta).$
+Proof.
+The proof can be found in Section B. ∎
+Remark 4.4.
+A careful review of the proof of Theorem 4.3 shows that, for this result, it suffices that the set $C$ be locally a convex body in a neighborhood of the (necessarily unique) projection point.
 The next example shows that the order of the bound obtained in Theorem 4.3 is sharp.
 Example 4.1.
-Consider the half-space $C=\{y\in\mathbb{R}^{n}:\langle y,n\rangle\leq 0\}$, where $\|\nu\|=1$.
+Consider the half-space $C=\{y\in\mathbb{R}^{n}:\langle y,\nu\rangle\leq 0\}$, where $\|\nu\|=1$.
 Then $\partial C=\{y\in\mathbb{R}^{n}:\langle y,\nu\rangle=0\}$ is a hyperplane.
@@
 $\operatorname{proj}_{C}(x)=x-\langle x,\nu\rangle\nu,\quad d_{C}(x)=\langle x,\nu\rangle.$
-Moreover, since $p_{\delta}(x)=\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}\!\left[\,y\mid y\in C\,\right]$, symmetry implies that the tangential conditional mean is unchanged and the drift is purely normal.
+Moreover, since $p_{\delta}(x)=\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}\!\left[\,y\mid y\in C\,\right]$, symmetry implies that the conditional mean in tangential directions is unchanged and the drift is purely normal.
 Hence,
@@
 $\alpha_{\delta}:=d_{C}(x)-\sqrt{\delta}\frac{\phi(d_{C}(x)/\sqrt{\delta})}{\Phi(-d_{C}(x)/\sqrt{\delta})},$
-and therefore
+where $\phi$ and $\Phi$ denote the standard normal density and distribution function, respectively. Therefore
 $p_{\delta}(x)=\operatorname{proj}_{C}(x)+\left(d_{C}(x)-\sqrt{\delta}\frac{\phi(d_{C}(x)/\sqrt{\delta})}{\Phi(-d_{C}(x)/\sqrt{\delta})}\right)\nu,$
 which implies that $\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|=O(\delta)$ as $\delta\to 0^{+}$. Hence, the order of the bound obtained in Theorem 4.3 is sharp.
+Acknowledgements
+D.
+Morales was supported by ANID Chile under the CMM BASAL funds for the Center of Excellence FB210005 and ANID BECAS/DOCTORADO NACIONAL 21252113.
+P.
+Pérez-Aros was supported by ANID (Chile) through Fondecyt Regular grants No.
+1220886, No.
+1240120, and No.
+1261728; CMM BASAL funds for the Center of Excellence FB210005; and the projects ECOS230027 and MATH-AmSud 23-MATH-17.
+E.
+Vilches was supported by ANID (Chile) through Fondecyt Regular grants No.
+1220886, No.
+1240120, and No.
+1261728; CMM BASAL funds for the Center of Excellence FB210005; and the projects ECOS230027 and MATH-AmSud 23-MATH-17.
 Appendix A Auxiliary lemmas
 Lemma A.1.
-Let $D\subset\mathbb{R}^{n}$ be a closed convex set with nonempty interior, and for $R>0$ set $\Omega_{R}:=\operatorname{int}(D)\cap\mathbb{B}_{R}$.
+Let $D\subset\mathbb{R}^{n}$ be a closed convex set with nonempty interior, and for $R>0$ set $\Omega_{R}:=\operatorname{int}(D)\cap B_{R}$.
 Then $\Omega_{R}$ is a bounded Lipschitz domain.
@@
 $R>0$,
-(8) $\mathcal{H}^{n-1}\big(\partial D\cap\partial\mathbb{B}_{R}\big)=0.$
+(6) $\mathcal{H}^{n-1}\big(\partial D\cap\partial B_{R}\big)=0.$
 Let $F\in W^{1,1}_{\operatorname{loc}}(\operatorname{int}(D);\mathbb{R}^{n})$. Then, for every $R>0$ such that (6) holds and $\int_{\Omega_{R}}(|F(x)|+\|\nabla F(x)\|)\,dx<\infty$, we have
-$\int_{\Omega_{R}}\operatorname{div}F(x)\,dx=\int_{\partial D\cap\mathbb{B}_{R}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\operatorname{int}(D)\cap\partial\mathbb{B}_{R}}F\cdot\nu_{\mathbb{B}_{R}}\,d\mathcal{H}^{n-1},$
+$\int_{\Omega_{R}}\operatorname{div}F(x)\,dx=\int_{\partial D\cap B_{R}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\operatorname{int}(D)\cap\partial B_{R}}F\cdot\nu_{B_{R}}\,d\mathcal{H}^{n-1},$
 where $\nu_{D}$ is the outer unit normal to $D$, defined $\mathcal{H}^{n-1}$-a.e.
-on $\partial D$, $\nu_{\mathbb{B}_{R}}$ is the outer unit normal to $\mathbb{B}_{R}$.
-Here, $F$ on $\partial\Omega_{R}$ denotes the interior Sobolev trace of $F|_{\Omega_{R}}$ on $\partial\Omega_{R}$, restricted to $\partial D\cap\mathbb{B}_{R}$ and to $\operatorname{int}(D)\cap\partial\mathbb{B}_{R}$.
+on $\partial D$, $\nu_{B_{R}}$ is the outer unit normal to $B_{R}$.
+Here, $F$ on $\partial\Omega_{R}$ denotes the interior Sobolev trace of $F|_{\Omega_{R}}$ on $\partial\Omega_{R}$, restricted to $\partial D\cap B_{R}$ and to $\operatorname{int}(D)\cap\partial B_{R}$.
 Proof.
@@
 $\mathcal{H}^{n-1}(\Phi_{k}(g_{k}^{-1}(R)))\leq\operatorname{Lip}(\Phi_{k})^{n-1}\mathcal{L}^{n-1}(g_{k}^{-1}(R))=0.$
-Moreover, $\Phi_{k}(g_{k}^{-1}(R))\subset\partial D\cap\partial\mathbb{B}_{R}$. Hence, for a.e. $R>0$,
-$\mathcal{H}^{n-1}\big((\partial D\cap\partial\mathbb{B}_{R})\cap\Phi_{k}(U_{k})\big)=0.$
+Moreover, $\partial D\cap\Phi_{k}(g_{k}^{-1}(R))\subset\partial D\cap\partial B_{R}$ and $\partial D\cap\partial{B}_{R^{\prime}}\cap\Phi_{k}(g_{k}^{-1}(R))=\emptyset$ for all $R^{\prime}\neq R$.
+Hence,
+$(\partial D\cap\partial B_{R})\cap\Phi_{k}(U_{k})=(\partial D\cap\partial B_{R})\cap\Phi_{k}(g^{-1}_{k}(R)),$
+which justifies that
+$\mathcal{H}^{n-1}\big((\partial D\cap\partial B_{R})\cap\Phi_{k}(U_{k})\big)=0.$
 Taking the countable union over $k$ and using the $\mathcal{H}^{n-1}$-negligibility of the uncovered part of $\partial D$, we obtain (6).
@@
 Next, we observe that
-$\partial\Omega_{R}=(\partial D\cap\mathbb{B}_{R})\cup(\operatorname{int}(D)\cap\partial\mathbb{B}_{R})\cup(\partial D\cap\partial\mathbb{B}_{R}).$
-Moreover, on $\partial D\cap\mathbb{B}_{R}$ the boundary $\partial\Omega_{R}$ coincides locally with $\partial D$.
+$\partial\Omega_{R}=(\partial D\cap B_{R})\cup(\operatorname{int}(D)\cap\partial B_{R})\cup(\partial D\cap\partial B_{R}).$
+Moreover, on $\partial D\cap B_{R}$ the boundary $\partial\Omega_{R}$ coincides locally with $\partial D$.
 Hence $\nu_{\Omega_{R}}=\nu_{D}$ $\mathcal{H}^{n-1}$-a.e.
-on $\partial D\cap\mathbb{B}_{R}$.
-Finally, by (8), the set $\partial D\cap\partial\mathbb{B}_{R}$ has $\mathcal{H}^{n-1}$-measure zero and does not contribute to the boundary integral in (9).
+on $\partial D\cap B_{R}$.
+Finally, by (6), the set $\partial D\cap\partial B_{R}$ has $\mathcal{H}^{n-1}$-measure zero and does not contribute to the boundary integral in (7).
 Therefore (7) splits into
-$\int_{\Omega_{R}}\operatorname{div}F\,dx=\int_{\partial D\cap\mathbb{B}_{R}}\operatorname{Tr}(F)\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\operatorname{int}(D)\cap\partial\mathbb{B}_{R}}\operatorname{Tr}(F)\cdot\nu_{\mathbb{B}_{R}}\,d\mathcal{H}^{n-1},$
+$\int_{\Omega_{R}}\operatorname{div}F\,dx=\int_{\partial D\cap B_{R}}\operatorname{Tr}(F)\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\operatorname{int}(D)\cap\partial B_{R}}\operatorname{Tr}(F)\cdot\nu_{B_{R}}\,d\mathcal{H}^{n-1},$
 as claimed. ∎
+The following lemma can be proved directly using polar coordinates and integration by parts.
 Lemma A.2.
@@
 $\int_{\{y\in\mathbb{R}^{n}:\ \|y\|\geq R\}}(\|y\|+d)^{k}\exp\!\Big(-\frac{\|y\|^{2}}{2\delta}\Big)\,dy\leq\frac{4\pi^{n/2}}{\Gamma\!\left(\frac{n}{2}\right)}(R+d)^{k}R^{n-2}\delta\,\exp\!\Big(-\frac{R^{2}}{2\delta}\Big).$
+The following lemma can be found in [6].
 Lemma A.3.
@@
 Moreover,
-$|\tilde{p}_{\delta}^{\nu}|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{2})\textrm{ as }\delta\to 0^{+}.$
+$|\tilde{p}_{\delta}^{\nu}|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{3/2})\textrm{ as }\delta\to 0^{+}.$
 Proof of Claim 8: Let $g_{\delta}(z)=\frac{1-r(A(z),\delta)}{A(z)}$. Then, by virtue of Claim 7, we get
@@
 Moreover, by Taylor’s theorem,
-$\frac{1}{\mathbb{E}_{\eta_{\delta}}[g_{\delta}(z)]}=\frac{1}{\frac{1}{d}+\mathbb{E}_{\eta_{\delta}}[g_{\delta}(z)-\frac{1}{d}]}=d-d^{2}\mathbb{E}_{\eta_{\delta}}[g_{\delta}(z)-\frac{1}{d}]+O(\delta^{2}).$
+$\frac{1}{\mathbb{E}_{\eta_{\delta}}[g_{\delta}(z)]}=\frac{1}{\frac{1}{d}+\mathbb{E}_{\eta_{\delta}}\left[g_{\delta}(z)-\frac{1}{d}\right]}=d-d^{2}\mathbb{E}_{\eta_{\delta}}\left[g_{\delta}(z)-\frac{1}{d}\right]+O(\delta^{2}).$
 Therefore,
 $\tilde{p}_{\delta}^{\nu}=\mathbb{E}_{\eta_{\delta}}[h(z)]-d\,\mathbb{E}_{\eta_{\delta}}[r(A(z),\delta)]+O(\delta^{2}).$
-Finally, using that $|h(z)|\leq\frac{L}{2}\|z\|^{2}$ and $\mathbb{E}_{\eta_{\delta}}[\|z\|^{2}]=(n-1)\delta+O(\delta^{2})$, we obtain that
-$|\tilde{p}_{\delta}^{\nu}|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{2}),$
+Finally, using that $|h(z)|\leq\frac{L}{2}\|z\|^{2}$ and $\mathbb{E}_{\eta_{\delta}}[\|z\|^{2}]\leq(n-1)\delta+O(\delta^{3/2})$, we obtain that
+$|\tilde{p}_{\delta}^{\nu}|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{3/2}),$
 which proves the claim.
@@
 By Claim 8 (together with Claims 1,5,6 to control the tail and the exponentially small terms), we obtain
-$|p_{\delta}^{\nu}(x)|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{2}).$
+$|p_{\delta}^{\nu}(x)|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{3/2}).$
 By Claim 10,